### v1

# Explainable Graph Neural Networks Under Fire

Abstract
Predictions made by graph neural networks (GNNs) usually lack interpretability due to their complex computational behavior and the abstract nature of graphs. In an attempt to tackle this, many GNN explanation methods have emerged. Their goal is to explain a model’s predictions and thereby obtain trust when GNN models are deployed in decision critical applications. Most GNN explanation methods work in a post-hoc manner and provide explanations in the form of a small subset of important edges and/or nodes. In this paper we demonstrate that these explanations can unfortunately not be trusted, as common GNN explanation methods turn out to be highly susceptible to adversarial perturbations. That is, even small perturbations of the original graph structure that preserve the model’s predictions may yield drastically different explanations. This calls into question the trustworthiness and practical utility of post-hoc explanation methods for GNNs. To be able to attack GNN explanation models, we devise a novel attack method dubbed GXAttack, the first optimization-based adversarial attack method for post-hoc GNN explanations under such settings. Due to the devastating effectiveness of our attack, we call for an adversarial evaluation of future GNN explainers to demonstrate their robustness.
1 Introduction
Graph Neural Networks (GNNs) have achieved promising results in various learning tasks, including representation learning, node classification, graph classification, link prediction, and anomaly detection (Zhou et al., 2020). However, intricate data representations and non-linear transformations render their interpretation and thus understanding their predictions non-trivial (Amara et al., 2022).
This has led to the introduction of GNN explanation methods, which can improve a model’s transparency and thus increase trust in a GNN model, especially when it is deployed in a decision-critical application (e.g., where fairness, privacy, or safety are important) (Ying et al., 2019). Moreover, they can help identify the scenarios where a GNN model may fail (Vu and Thai, 2020). Explainability of a GNN model can be obtained by designing a self-explainable GNN model, which usually employs a simple model architecture and few parameters, leading to explainability at the cost of suboptimal prediction accuracy; or by extracting post-hoc explanations that do not influence the internal workings of a GNN model and thus maintains its high prediction accuracy. In this paper we only consider post-hoc GNN explanation methods, as many such methods have merged over the past decade, including but not limited to Ying et al. (2019); Pope et al. (2019); Yuan et al. (2020); Luo et al. (2020); Vu and Thai (2020); Bajaj et al. (2021); Yuan et al. (2021); Lucic et al. (2022); Huang et al. (2022).
It has been shown that traditional deep neural networks (DNNs) are vulnerable to adversarial attacks (Szegedy et al., 2013; Chakraborty et al., 2018), where an imperceptible well-designed perturbation to input data can lead to substantially different prediction results. Recently, it was demonstrated that the explanations of predictions made by these DNNs are also fragile (Baniecki and Biecek, 2024). That is, an unnoticeable, well-designed perturbation to input data can result in completely different explanations while the model’s predictions remain unchanged. However, the vulnerability of explanation methods has been primarily studied for DNN models designed for image and text data. In contrast, the vulnerability of GNN explanation methods to adversarial attacks has received very limited attention (Fang et al., 2024; Wang et al., 2024), although a plethora of GNN explanation methods have recently been proposed (Yuan et al., 2022; Li et al., 2022; Kakkad et al., 2023).
Importantly, existing studies have only investigated the (in)stability of GNN explanations under relatively “mild” settings: they only consider random perturbations and/or prediction-altering perturbations. Specifically, Agarwal et al. (2022) theoretically analyzed the stability of three GNN explainers under random perturbations. Further, Wang et al. (2024) considered two scenarios: 1) graphs whose predictions are not changed after randomly flipping $20\%$ edges; and 2) severe perturbations using an off-the-shelf attack algorithm to perturb at most $10\%$ edges to change the GNN predictions (as well as the explanations). Besides, Fan et al. (2023) proposed an adversarial attack to simultaneously change the GNN predictions and their explanations. By altering predictions and explanations, however, the graph can become inherently different. Fang et al. (2024) explored evaluation metrics for GNN explanations from the perspective of adversarial robustness and resistance to out-of-distribution input, where they defined “adversarial robustness" as the difficulty of reversing a GNN prediction by perturbing the complementary of the explanatory subgraph.
Figure 1: When using post-hoc GNN explainers, the explanatory subgraph on a graph with “prediction-preserving" perturbations (right) can strongly differ from that on the original graph (left).
Different from previous studies, we aim to perform perturbations that are both carefully crafted (i.e., optimized rather than random) and imperceptible (i.e., such that the GNN predictions remain unchanged but their explanations differ substantially). As shown in Figure 1, we empirically found that small prediction-preserving perturbations can result in largely different explanations generated by post-hoc GNN explainers. To our knowledge, we are the first to formally formulate and systematically investigate this problem. Specifically, we devise GXAttack, the first optimization-based adversarial attack on post-hoc GNN explanations under this setting. We employ a widely used GNN explainer, PGExplainer (Luo et al., 2020), as example target when designing our attack algorithm. Results on various datasets demonstrate the effectiveness of our approach. Moreover, our experiments show that other widely used GNN explanation methods, such as GradCAM (Pope et al., 2019), GNNExplainer (Ying et al., 2019), and SubgraphX (Yuan et al., 2021), are also fragile under the attacks optimized for PGExplainer. Due to the devastating effectiveness of our attack, we call for an adversarial evaluation of future GNN explainers to demonstrate their robustness.
2 Preliminaries
We utilize lowercase letters, bold lowercase letters, uppercase letters and calligraphic fonts to represent scalars ($x$), vectors ($\mathbf{x}$), matrices ($\mathbf{X}$), and sets ($\mathcal{X}$), respectively.
Definition 1 (Attributed Graph).
We denote an attributed graph as $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{X}\}$ where $\mathcal{V}=\{v_{1},...,v_{n}\}$ is the set of nodes. Let =$\mathcal{E}=\{e_{ij}\}_{i,j\in\{1,...,n\}}$ be the set of edges, where $e_{ij}=1$ if there exists an edge between $v_{i}$ and $v_{j}$ and $e_{ij}=0$ otherwise. We use $\mathbf{A}\in\{0,1\}^{n\times n}$ to denote the corresponding adjacency matrix. $\mathbf{X}\in\mathbb{R}^{n\times d}$ represents the node attribute matrix, where the $i$-th row vector $\mathbf{x}_{i}$ denotes the node attribute of $v_{i}$. Thus, a graph can also be represented as $\mathcal{G}=(\mathbf{A},\mathbf{X})$.
2.1 Node classification with GNNs
We consider the task of node classification. Given an input graph $\mathcal{G}=\{\mathcal{V},\mathcal{E},\mathbf{X}\}$ (or $\mathcal{G}=(\mathbf{A},\mathbf{X})$), a prediction model $f_{\theta}(\cdot)$ assigns a label $c$ from a pre-defined set $\mathcal{C}$ to each node $v\in\mathcal{V}$. In other words, we have $f_{\theta}(\mathcal{G}(v))=c$, where $\mathcal{G}(v)$ denotes the computation graph of node $v$. Typically, $f_{\theta}(\cdot)$ consists of 1) a GNN model used to learn node representations, and 2) a multi-layer perceptron (MLP) classifier that assigns labels. More specifically, given a set of training nodes and their labels $\{(v_{1},y_{1}),...,(v_{n},y_{n})\}$ from graph $\mathcal{G}$, the objective function of a GNN classifier can be defined as
[…]
where $\theta$ is the set of learnable parameters for $f_{\theta}(\cdot)$, $l(\cdot)$ is the cross-entropy loss function, and $\mathbb{I}(\cdot)$ denotes the Kronecker function. Further, $f_{\theta}(\mathcal{G}(v_{i}),y_{i})^{(c)}$ indicates the $c$-th softmax output for the prediction of node $v_{i}$, and $C$ is the cardinality of label set $\mathcal{C}$.
Now we define the design detail of the GNN classifier $f_{\theta}(\cdot)$. For the first component, for simplicity, we employ a 2-layer GCN model with parameters $(\mathbf{W}_{1},\mathbf{W}_{2})$, i.e.,
$GCN(\mathcal{G})=GCN(\mathbf{A},\mathbf{X}):=\tilde{\mathbf{A}}\sigma(\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}_{1})\mathbf{W}_{2},$ (2)
where $\tilde{\mathbf{A}}=\tilde{\mathbf{D}}^{-1/2}(\mathbf{A}+\mathbf{I})\tilde{\mathbf{D}}^{-1/2}$ (with $\mathbf{I}$ the identity matrix and $\tilde{\mathbf{D}}$ the diagonal matrix of $\mathbf{A}+\mathbf{I}$). Further, $\sigma(\cdot)$ is an activation function such as ReLU. For the second component, we simply use a $\operatorname{softmax}(\cdot)$ function rather than a MLP. Therefore, the full GNN classifier $f_{\theta}$ is defined as
$f_{\theta}(\mathbf{A},\mathbf{X}):=\operatorname{softmax}(\tilde{\mathbf{A}}\sigma(\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}_{1})\mathbf{W}_{2}).$ (3)
2.2 GNN Explainer
We consider post-hoc GNN explanation methods that provide instance-level explanations: given a graph $\mathcal{G}$, a target node $v$, and prediction $f_{\theta}(\mathcal{G}(v))$, the explainer $g_{\lambda}(v,f(\mathcal{G}(v)))$ aims to ‘explain’ this particular prediction. Following common explanation methods, we assume that explanations are given in the form of an explanatory subgraph, namely $\mathcal{G}_{s}(v)=g(v,f(\mathcal{G}(v)))$, consisting of a small set of important edges and their associated nodes, where the explanatory subgraph is usually connected and edge importance is measured by a so-called edge importance score.
In this paper, we take PGExplainer (Luo et al., 2020) as an example to attack for three reasons. First, Faber et al. (2021) pointed out that GNN explanations providing edge importance are preferable to explanations providing node importance, due to the fact that edges have more fine-grained information than nodes. Second, Agarwal et al. (2023) empirically showed that PGExplainer generates the least unstable explanations, exhibiting a $35.35\%$ reduction in explanation instability relative to the average instability of other GNN explainers. Third, the objective function of PGExplainer is differentiable.
For completeness, we briefly revisit the goal of PGExplainer (see Appendix A for more details). For brevity, we denote $\mathcal{G}(v)$ as $\mathcal{G}$ (and $\mathcal{G}_{s}(v)$ as $\mathcal{G}_{s}$ analogously) throughout the remainder of the paper.
PGExplainer. Given a graph $\mathcal{G}$ and a target node $v$, PGExplainer provides an explanatory subgraph $\mathcal{G}_{s}(v)=(\mathbf{A}_{s},\mathbf{X}_{s})$ to explain why $f_{\theta}(\mathcal{G}(v))=c$, where $\mathbf{A}_{s}$ is the subgraph structure and $\mathbf{X}_{s}$ is the node features associated with nodes residing in the subgraph. Specifically, PGExplainer maximizes the mutual information between the GNN’s prediction $\mathbf{y}$ and the explanatory subgraph $\mathcal{G}_{s}$, i.e.,
$\underset{\mathcal{G}_{s}}{\max}~{}\text{MI}(\mathbf{y},\mathcal{G}_{s}):=H(\mathbf{y})-H(\mathbf{y}|\mathcal{G}_{s}),$ (4)
where $H(\mathbf{y})$ is the entropy of $f_{\theta}(\mathcal{G})$, and $H(\mathbf{y}|\mathcal{G}_{s})$ is the conditional entropy of the GNN’s prediction on the explanatory graph, namely $f_{\theta}(\mathcal{G}_{s})$.
3 GXAttack: GNN Explanation Adversarial Attacks
Given an input $\mathcal{G}$, a GNN classifier $f_{\theta}(\cdot)$, and a GNN explainer $g_{\lambda}(\cdot)$, the original prediction for the target node $v$ is $f_{\theta}(\mathcal{G}(v))$ and its corresponding original explanatory subgraph is given by $\mathcal{G}_{s}(v)=g_{\lambda}(f_{\theta}(v,\mathcal{G}(v)))$. From an attacker’s perspective, a desirable manipulated graph $\mathcal{G}_{adv}:=\hat{\mathcal{G}}=\phi({\mathcal{G}})$ for attacking node $v$ should have the following properties, where $\phi(\cdot)$ is an attack function that perturbs its input, i.e., it removes and/or adds a small set of edges:
  1. 1.
    The norm of the perturbation is so small that the change is considered imperceptible, i.e., $\mathcal{G}\approx\hat{\mathcal{G}}$. This can be formalized as $\|\mathbf{A}-\hat{\mathbf{A}}\|<B$, where $B$ is the perturbation budget;
  2. 2.
    As GNNs can be highly sensitive to input perturbations, we require that the output of the GNN classifier remains approximately the same, i.e., perturbations must be prediction-preserving. Formally, we have $f(\mathcal{G}(v))=f(\hat{\mathcal{G}}(v))$. We hypothesize that perturbations that are both small and prediction-preserving are very likely to preserve the essence of the learned predictive models;
  3. 3.
    The explanatory subgraph for node $v$ on original graph $\mathcal{G}_{s}(v)=g_{\lambda}(f_{\theta}(v,\mathcal{G}(v)))$ and that on perturbed graph $\hat{\mathcal{G}}_{s}(v)=g_{\lambda}(f_{\theta}(\hat{\mathcal{G}}(v)))$ are substantially different: $\mathcal{G}_{s}(v)\neq\hat{\mathcal{G}}_{s}(v)$. In particular, we aim to make $\hat{\mathbf{A}}_{s}$ strongly different from $\mathbf{A}_{s}$.
3.1 Attack Objectives
Generic Objective. This paper focuses on adversarial structure attacks on post-hoc GNN explanation for node classification:
[…]
with loss function $\mathcal{L}$, perturbation budget $B$, and fixed parameters $\theta$ and $\lambda$. The GNN predictor $f_{\theta}(\cdot)$ is applied to a graph $\mathcal{G}=\{\mathbf{A},\mathbf{X}\}$ to obtain node label predictions $f_{\theta}(\mathbf{A},\mathbf{X})$. Next, GNN explainer $g_{\lambda}(\cdot)$ is applied to GNN predictions $f_{\theta}(\mathbf{A},\mathbf{X})$ to generate post-hoc explanations $g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X}))$. We focus on evasion (test time) attacks and local attacks on a single node with a budget $B$. Moreover, we study white box attacks as they represent the “worst-case noise” of a model. In other words, we assume perfect knowledge about the graph, labels, prediction model, and post-hoc explanation model.
Loss Function. We should instantiate the loss function $\mathcal{L}$ such that it pertains to the desirable properties of manipulations discussed before. More specifically, it is defined as
$\mathcal{L}:=\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))).$ (6)
The first term ensures that the GNN predictions before and after perturbations are approximately the same. Further, the second term enforces that the manipulated explanation is different from the original explanatory subgraph. Their relative importance is weighted by hyperparameter $\beta\in\mathbb{R}_{+}$.
Final Objective. We directly model the edge perturbation matrix $\mathbf{P}\in\{0,1\}^{n\times n}$, where we flip $\mathbf{A}_{ij}$ (namely $0\to 1$ or $1\to 0$) if $\mathbf{P}_{ij}=1$, and keep $\mathbf{A}_{ij}$ intact otherwise. On this basis, we have
$\underset{\mathbf{P}\in\{0,1\}^{n\times n}\text{~{}s.t.~{}}\sum\mathbf{P}_{ij}<B}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))),$ (7)
where $\oplus$ means element-wise exclusive or operation, and $B$ is the edge flipping budget.
3.2 Relaxations and Loss Definitions
Unfortunately, directly minimizing $\mathcal{L}$ w.r.t. $\mathbf{P}$ is intractable since there are $\mathcal{O}(2^{n\times n})$ candidates for $\mathbf{P}$. We therefore approximate this discrete optimization problem, i.e., we make some relaxations to make $\mathcal{L}$ (approximately) differentiable w.r.t. $\mathbf{P}$ so that we can perform gradient-based attacks.
Relaxations. We assume that edge flipping graph $\mathcal{M}$ corresponding to $\mathbf{P}\in\{0,1\}^{n\times n}$ (namely using $\mathbf{P}$ as the adjacency matrix) is a Gilbert random graph. In other words, edge occurrences are conditionally independent from each other. As a result, with slight abuse of notations, the probability of graph $\mathcal{M}$ can be factorized as $P(\mathcal{M})=\prod_{1\leq i,j\leq n}P(m_{ij})$, dubbed $\tilde{\mathbf{P}}$. Next, we instantiate $P(m_{ij})$ using a Bernoulli distribution, namely $P(m_{ij})\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{ij})$ with $P(m_{ij}=1)=\tilde{\mathbf{P}}_{ij}$. As a result, Eq 7 is equivalent to
$\underset{\mathbf{P}}{{\max}}~{}\mathcal{L}(\mathbf{P})=\underset{\mathcal{M}}{\max}~{}\mathbb{E}_{\mathcal{M}}[\mathcal{L}(\mathcal{M})]\approx\underset{\tilde{\mathbf{P}}}{\max}~{}\mathbb{E}_{\mathcal{M}\sim\tilde{\mathbf{P}}}[\mathcal{L}(\mathcal{M})].$ (8)
In this way, we relax the $\mathbf{P}\in\{0,1\}^{n\times n}$ from binary variables to continuous variables $\tilde{\mathbf{P}}\in[0,1]^{n\times n}$. Hence, we can utilise gradient-based methods to optimize the objective function. At the last step of the attack, we then discretize the result (see Section 3.3).
Loss Term 1. We approximate the first term in $\mathcal{L}$ as follows:
$\underset{\tilde{\mathbf{P}}}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X})):=\underset{\tilde{\mathbf{P}}}{\max}~{}H(f_{\theta}(\mathbf{A},\mathbf{X}))-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))\$ (9)
$\propto\underset{\tilde{\mathbf{P}}}{\max}~{}-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))\approx\underset{\tilde{\mathbf{P}}}{\max}~{}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}})).$ (10)
Eq 9 holds by definition of mutual information; Eq 10 left half holds as $\theta,\mathbf{A},\mathbf{X}$ are fixed; Eq 10 right half holds when we replace conditional entropy $H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ with cross entropy $H(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$. We define $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}+\tilde{\mathbf{P}}_{ij}$ if $\mathbf{A}_{ij}=0$, and $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}-\tilde{\mathbf{P}}_{ij}$ otherwise.
Loss Term 2. Next, we approximate the second term in $\mathcal{L}$ as follows:
$\underset{\tilde{\mathbf{P}}}{\min}~{}\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X})))$ $\propto\underset{\tilde{\mathbf{P}}}{\max}~{}\text{Dist}(\mathbf{M}_{s},\hat{\mathbf{M}}_{s})$ (11)
$=:\underset{\tilde{\mathbf{P}}}{\max}~{}\left(1-\frac{\text{vec}(\mathbf{M}_{pr})\cdot\text{vec}(\hat{\mathbf{M}}_{pr})}{\|\mathbf{M}_{pr}\|_{F}\|\hat{\mathbf{M}}_{pr}\|_{F}}\right),$ (12)
Eq 11 holds by definition; Eq 12 holds when we define the $Dist$ function as the cosine distance metric, where $\text{vec}(\cdot)$ flattens a matrix to a vector, and $\|\cdot\|_{F}$ represents the Frobenius Norm.
Final Relaxed Objective. As a result, optimizing Eq 7 amounts to optimizing
$\underset{\tilde{\mathbf{P}}\in[0,1]^{n\times n}\text{~{}s.t.~{}}\sum\tilde{\mathbf{P}}_{ij}<B}{\max}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}}))+\beta*\left(1-\frac{\text{vec}(\mathbf{M}_{pr})\cdot\text{vec}(\hat{\mathbf{M}}_{pr})}{\|\mathbf{M}_{pr}\|_{F}\|\hat{\mathbf{M}}_{pr}\|_{F}}\right),$ (13)
where $\mathcal{L}(\tilde{\mathbf{P}})$ is differentiable w.r.t. $\tilde{\mathbf{P}}$.
3.3 Optimization
We employ Projected Gradient Ascent Algorithm (based on Xu et al. (2019)) to optimize objective function (13), and the details are given in Algorithm 1. To go back from the continuous $\tilde{\mathbf{P}}_{ij}$ to discrete ${\mathbf{P}}_{ij}$, we perform Bernoulli sampling with ${\mathbf{P}}_{ij}\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{ij})$ after optimization (Line 10). We note that the Projected Randomized Block Coordinate Descent (PR-BCD) (Geisler et al., 2021) can perform this optimization more efficiently, but do not compare both methods due to the low scalability of the evaluated explainers. The complexity analysis of GXAttack is given in Appendix C.
Algorithm 1 Pseudo-code for GXAttack
1:graph $\mathcal{G}=(\mathbf{A},\mathbf{X})$, predictor $f_{\theta}(\cdot)$, explainer $g_{\lambda}(\cdot)$, loss $\mathcal{L}$, budget $B$, #epochs $T$, learning rate $\alpha_{t}$
2:perturbed adjacency matrix $\hat{\mathbf{A}}$
3:$\mathbf{y}\leftarrow f_{\theta}(\mathbf{A},\mathbf{X})$triangleright Get predictions on original graph
[…]
5:$\tilde{\mathbf{P}}\leftarrow\mathbf{0}\in\mathbb{R}^{n\times n}$triangleright Initialize zeros for continuous perturbation variables
6:for $t\in\{1,2,...,T\}$ do
7:    $\hat{\mathbf{y}}\leftarrow f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}}_{t-1},\mathbf{X})$triangleright Get predictions on perturbed graph
8:    $\hat{\mathcal{G}}_{s}\leftarrow g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}}_{t-1},\mathbf{X})$$\triangleright$ Get explanations for the predictions on perturbed graph
9:    $\tilde{\mathbf{P}}_{t}\leftarrow\tilde{\mathbf{P}}_{t-1}+\alpha_{t}\cdot\nabla_{\tilde{\mathbf{P}}_{t-1}}\mathcal{L}(\hat{\mathbf{y}},\mathbf{y},\hat{\mathcal{G}}_{s},\mathcal{G}_{s})$triangleright Update gradients
10:    $\tilde{\mathbf{P}}_{t}\leftarrow\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})\leq B]}(\tilde{\mathbf{P}}_{t})$triangleright Project gradients, see Appendix B for more detail
11:end for
12:$\mathbf{P}\sim\text{Bernoulli}(\tilde{\mathbf{P}})$ s.t. $\sum\mathbf{P}_{ij}\leq B$triangleright Sample binary perturbation variables
13:Return $\mathbf{A}\oplus\mathbf{P}$triangleright Which is defined as $\mathbf{A}_{ij}+\mathbf{P}_{ij}(1-2\mathbf{A}_{ij})$
4 Experimental Setup
We aim to answer the following research questions: (RQ1) How effective are the attacks generated by GXAttack on PGExplainer? (RQ2) How effective are the attacks generated by GXAttack when compared to trivial attackers that use random perturbations? and (RQ3) Can the attacks generated by GXAttack (optimized for PGExplainer) transfer to other GNN explanation methods?
4.1 Datasets
To answer these questions, we need to evaluate the GNN explanation quality before and after the attacks. However, as pointed out by Agarwal et al. (2023), evaluating the quality of post-hoc GNN explanations is challenging. This is because existing benchmark graph datasets lack ground-truth explanations or their explanations are not reliable. To mitigate this, they proposed a synthetic graph data generator ShapeGGen. It can generate a variety of graph datasets with ground-truth explanations, where graph size, node degree distributions, the ground-truth explanation sizes, and more can be varied. As a result, ShapeGGen allows us to mimic graph data in various real-world applications. More importantly, ShapeGGen can ensure that the generated data and its ground-truth explanations do not suffer from GNN explanation evaluation pitfalls (Faber et al., 2021), namely trivial explanations, redundant explanations, and weak GNN predictors. This makes ShapeGGen very suitable for studying the limitations of GNN explainers. We consider the synthetic datasets in Table 3 (See Appendix D).
Why not real-world datasets? We do not consider real-world datasets for two key reasons: 1) synthetic datasets with ground truth are de-facto standards for evaluating post-hoc GNN explainers (for node classification), including GNNExplainer, PGExplainer, PGM-Explainer, SubgraphX, CF-GNNExplainer, RCExplainer, etc. (see Table 11 for more details); 2) without ground-truth information we can only compute some metrics, such as cosine similarity of explanation before and after attack (which will be defined later) to quantify attack performance. However, as shown in Table 1, the cosine similarity and the real attack performance (in terms of $\Delta\text{GEA}$, which will be defined later) are not necessarily related. Therefore, considering such metrics may not be meaningful.
4.2 Metrics and Baselines
Graph Explanation Accuracy (GEA). We employ the graph explanation accuracy proposed in Agarwal et al. (2023) to quantify the quality of generated explanations when the ground-truth explanations are available. For simplicity, we assume that each edge is binary coded as ‘0’ (non-contributory) or ‘1’ (contributory) to model predictions in both ground-truth (provided by ShapeGGen) and predicted explanation masks (provided by the GNN explainer). Specifically, graph explanation accuracy measures the correctness of generated explanations by computing the Jaccard Index between ground-truth explanation $\mathbf{M}_{gt}$ and predicted explanation $\mathbf{M}_{pr}$ as
$\text{JAC}(\mathbf{M}_{gt},\mathbf{M}_{pr})=\frac{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})}{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FN}(\mathbf{M}_{gt},\mathbf{M}_{pr})},$ (14)
where TP, FP, FN denote True Positives, False Positives, and False Negatives, respectively. Higher values indicate larger consistency between the generated explanation and the ground-truth explanation. As a result, the difference in this metric before and after perturbation can be used to measure the performance of attacks:
$\Delta\text{GEA}=:\text{JAC}(\mathbf{M}_{gt},\mathbf{M}_{pr})-\text{JAC}(\mathbf{M}_{gt},\hat{\mathbf{M}}_{pr}),$ (15)
where higher values indicates more successful attacks.
Cosine Similarity. To quantify the performance of attacks when ground-truth explanations are not available, we use cosine similarity to measure the stability of graph explanations:
$\operatorname{Sim}_{\cos}(\mathbf{M}_{pr},\hat{\mathbf{M}}_{pr})=\frac{\text{vec}(\mathbf{M}_{pr})\cdot\text{vec}(\hat{\mathbf{M}}_{pr})}{\|\mathbf{M}_{pr}\|_{F}\|\hat{\mathbf{M}}_{pr}\|_{F}},$ (16)
where $\text{vec}(\cdot)$ flattens its input as a vector, and $\|\cdot\|_{F}$ represents the Frobenius Norm. Higher values of cos_sim (within $[-1,1]$) indicate higher graph explanation stability, and thus less successful attacks.
Prediction Change. To measure the impact of edge perturbations on GNN predictions, we consider 1) $\Delta$Label: the ratio of nodes of which the predicted label has changed after attacks; and 2) $\Delta$Prob: the average absolute change of predicted probability (of the original predicted class) of all nodes.
Baselines. Given the novelty of our problem setting, we only consider two trivial baselines for edge perturbations: 1) random flipping: we randomly flip $B_{1}$ edges; and 2) random rewiring: we randomly rewire $B_{2}$ edges within k-hop neighbours of the target node.
4.3 Attack Transferability Settings
We consider attack transferability on the following four types of GNN explainers: 1) Gradient-based: Grad (Simonyan et al., 2013), GradCAM (Pope et al., 2019), GuidedBP (Baldassarre and Azizpour, 2019), Integrated Gradients (IG) (Sundararajan et al., 2017); 2) Perturbation-based: GNNExplainer (Ying et al., 2019), SubgraphX (Yuan et al., 2021); 3) Surrogate-based: PGMExplainer (Vu and Thai, 2020); and 4) Random Explainer that randomly generates random important edges.
For space reasons, all training protocols are deferred to Appendix G. Code is given in anonymous2024code.
5 Experimental Results and Analysis
Before answering the research questions, we state three presumptions and check them empirically in Section 5.1. Next, we answer RQ1 and RQ2 in Section 5.2, and RQ3 in Section 5.3.
5.1 Experiments to Check Presumptions
Presumption 1. The prediction correctness of a GNN predictor and explanation accuracy of a GNN explainer are not related. For each dataset, we first get the set of correctly predicted nodes and the set of wrongly predicted nodes; then we compare the distribution of their explanation accuracy. As shown in Figure 2 in Appendix E, there is no notable difference. In other words, we do not observe that wrongly labelled nodes tend to have lower explanation accuracy despite Faber et al. (2021) claim that it is unfair to use wrongly labelled nodes to evaluate GNN explanations (as they may not use the ground-truth explanations to make predictions and thus tend to have lower explanation accuracy).
Presumption 2. The prediction confidence of a GNN predictor and explanation accuracy of a GNN explainer are related, where prediction confidence is defined as the probability assigned to the predicted class. For each dataset, we first divide the nodes into five groups based on their prediction confidence; then we compare the distribution of their explanation accuracy. Tables 410 in Appendix E show that group ‘G9’ (i.e., the set of nodes with prediction confidence in $[0.9,1.0]$) indeed tends to have lower explanation accuracy on the original graph than other groups.
Presumption 3. The prediction confidence of a GNN predictor and attack performance of a GNN explanation attacker are related. To check this presumption, we first divide the nodes into five groups based on their prediction confidence; then we compare the distribution of their attack performance. Tables 410 in Appendix E show that group ‘G9’ also tends to have less successful attacks (i.e., smaller $\Delta\text{GEA}$) when compared to other groups. Possible reasons are that: 1) the nodes in this group have lower explanation accuracy on original graph; and 2) the prediction confidence is close to 1, leaving no much room for perturbations. These two facts together make the attack more challenging.
5.2 Experiments to Answer RQ1 and RQ2: Attack Effectiveness
Table 1: Results on all synthetic datasets (average $\pm$ standard deviations across 5 trials). O. GEA means explanation accuracy on original graph and P. GEA indicates explanation accuracy on perturbed graph. Moreover, $\Delta$GEA is the explanation accuracy change after perturbation, Simcos is the cosine similarity between explanations (before and after perturbations), and #Pert. represents the number of flipped edges. $\Delta$Label is the average change of predicted labels, while $\Delta$Prob denotes the average absolute change of predicted probability (of the original predicted class). “GXAttack", “Rnd. Flipping", and “Rnd. Rewiring" correspond to our attack method, random flipping, and random rewiring, respectively. $\uparrow^{\prime}$ indicates larger values are preferred while $\downarrow^{\prime}$ means the opposite.
Dataset Syn1 Syn2 Syn3 Syn4 Syn5 Syn6 Syn7
O. GEA $0.678_{\pm\text{0.046}}$ $0.641_{\pm\text{0.022}}$ $0.494_{\pm\text{0.012}}$ $0.469_{\pm\text{0.006}}$ $0.439_{\pm\text{0.014}}$ $0.306_{\pm\text{0.025}}$ $0.429_{\pm\text{0.013}}$
GXAttack P. GEA()($\downarrow$)( ↓ ) $0.497_{\pm\text{0.041}}$ $0.375_{\pm\text{0.011}}$ $0.308_{\pm\text{0.010}}$ $0.312_{\pm\text{0.003}}$ $0.290_{\pm\text{0.004}}$ $0.217_{\pm\text{0.013}}$ $0.287_{\pm\text{0.014}}$
$\Delta$GEA ()($\uparrow$)( ↑ ) $0.181_{\pm\text{0.020}}$ $0.267_{\pm\text{0.014}}$ $0.186_{\pm\text{0.013}}$ $0.157_{\pm\text{0.010}}$ $0.149_{\pm\text{0.015}}$ $0.088_{\pm\text{0.014}}$ $0.142_{\pm\text{0.009}}$
[…]
#Pert. ($\downarrow$) $3.300_{\pm\text{0.255}}$ $5.620_{\pm\text{0.045}}$ $6.680_{\pm\text{0.148}}$ $5.900_{\pm\text{0.122}}$ $6.560_{\pm\text{0.055}}$ $5.940_{\pm\text{0.167}}$ $7.020_{\pm\text{0.130}}$
$\Delta$Label($\downarrow$) $0.000_{\pm\text{0.000}}$ $0.002_{\pm\text{0.003}}$ $0.001_{\pm\text{0.001}}$ $0.001_{\pm\text{0.001}}$ $0.001_{\pm\text{0.001}}$ $0.000_{\pm\text{0.001}}$ $0.001_{\pm\text{0.001}}$
$\Delta$Prob($\downarrow$) $0.106_{\pm\text{0.008}}$ $0.140_{\pm\text{0.004}}$ $0.163_{\pm\text{0.003}}$ $0.150_{\pm\text{0.002}}$ $0.143_{\pm\text{0.004}}$ $0.089_{\pm\text{0.003}}$ $0.147_{\pm\text{0.001}}$
Rnd. Flipping P. GEA()($\downarrow$)( ↓ ) $0.640_{\pm\text{0.056}}$ $0.638_{\pm\text{0.023}}$ $0.493_{\pm\text{0.012}}$ $0.467_{\pm\text{0.009}}$ $0.438_{\pm\text{0.013}}$ $0.306_{\pm\text{0.024}}$ $0.429_{\pm\text{0.013}}$
$\Delta$GEA()($\uparrow$)( ↑ ) $0.038_{\pm\text{0.029}}$ $0.001_{\pm\text{0.005}}$ $0.001_{\pm\text{0.002}}$ $0.002_{\pm\text{0.003}}$ $0.001_{\pm\text{0.001}}$ $0.000_{\pm\text{0.005}}$ $0.000_{\pm\text{0.001}}$
[…]
#Pert.()($\downarrow$)( ↓ ) $15_{\pm\text{0.000}}$ $15_{\pm\text{0.000}}$ $15_{\pm\text{0.000}}$ $15_{\pm\text{0.000}}$ $15_{\pm\text{0.000}}$ $15_{\pm\text{0.000}}$ $15_{\pm\text{0.000}}$
$\Delta$Label()($\downarrow$)( ↓ ) $0.051_{\pm\text{0.012}}$ $0.011_{\pm\text{0.007}}$ $0.012_{\pm\text{0.004}}$ $0.009_{\pm\text{0.006}}$ $0.004_{\pm\text{0.002}}$ $0.002_{\pm\text{0.002}}$ $0.004_{\pm\text{0.002}}$
$\Delta$Prob()($\downarrow$)( ↓ ) $0.048_{\pm\text{0.003}}$ $0.009_{\pm\text{0.005}}$ $0.006_{\pm\text{0.001}}$ $0.008_{\pm\text{0.001}}$ $0.006_{\pm\text{0.001}}$ $0.003_{\pm\text{0.001}}$ $0.003_{\pm\text{0.001}}$
Rnd. Rewiring P. GEA()($\downarrow$)( ↓ ) $0.149_{\pm\text{0.012}}$ $0.045_{\pm\text{0.005}}$ $0.069_{\pm\text{0.004}}$ $0.067_{\pm\text{0.004}}$ $0.020_{\pm\text{0.002}}$ $0.013_{\pm\text{0.001}}$ $0.020_{\pm\text{0.002}}$
$\Delta$GEA()($\uparrow$)( ↑ ) $0.530_{\pm\text{0.043}}$ $0.596_{\pm\text{0.022}}$ $0.425_{\pm\text{0.010}}$ $0.402_{\pm\text{0.003}}$ $0.419_{\pm\text{0.014}}$ $0.293_{\pm\text{0.025}}$ $0.409_{\pm\text{0.012}}$
[…]
#Pert.()($\downarrow$)( ↓ ) $16_{\pm\text{0.000}}$ $16_{\pm\text{0.000}}$ $16_{\pm\text{0.000}}$ $16_{\pm\text{0.000}}$ $16_{\pm\text{0.000}}$ $16_{\pm\text{0.000}}$ $16_{\pm\text{0.000}}$
$\Delta$Label()($\downarrow$)( ↓ ) $0.274_{\pm\text{0.030}}$ $0.199_{\pm\text{0.022}}$ $0.202_{\pm\text{0.004}}$ $0.184_{\pm\text{0.011}}$ $0.214_{\pm\text{0.007}}$ $0.159_{\pm\text{0.013}}$ $0.274_{\pm\text{0.014}}$
$\Delta$Prob()($\downarrow$)( ↓ ) $0.234_{\pm\text{0.016}}$ $0.210_{\pm\text{0.013}}$ $0.198_{\pm\text{0.007}}$ $0.200_{\pm\text{0.003}}$ $0.221_{\pm\text{0.005}}$ $0.174_{\pm\text{0.013}}$ $0.259_{\pm\text{0.005}}$
From Table 1, we have the following main observations:
Observation 1. The larger the subgraph size, the lower the original explanation accuracy of PGExplainer and thus the poorer the attack performance (i.e., the smaller the value of $\Delta\text{GEA}$). For this, compare the results on Syn2 (Subgraph Size = 6) to those on Syn3 (Subgraph Size = 10).
Observation 2. The type of motifs (namely ground-truth explanations) appears to have minimal impact on both explanation accuracy and attack performance. This can be obtained by comparing the results on Syn3 (Motif =“House") with those on Syn4 (Motif =“Circle").
Observation 3. The larger the connection probability (namely node degree), the lower the original explanation accuracy of PGExplainer and thus the poorer the attack performance (i.e., the smaller the value of $\Delta\text{GEA}$). This can be obtained by comparing the results on Syn3 (P(Connection) = 0.06), Syn5 (P(Connection) = 0.12), and Syn6 (P(Connection) = 0.20).
Observation 4. The more subgraphs, the lower the original explanation accuracy of PGExplainer and thus the poorer the attack performance (i.e., the smaller the value of $\Delta\text{GEA})$. This can be obtained by comparing the results on Syn3 (#Subgraphs=50) with those on Syn7 (#Subgraphs=100).
Answer to RQ1. From Table 1, we can see that GXAttack can achieve a $\Delta\text{GEA}$ ranging from $0.088$ to $0.267$ while using a very small perturbation budget (less than 7.1 on all datasets) and keeping the predicted labels nearly unchanged (i.e., $\Delta\text{Label}$ is nearly zero) after perturbation.
Answer to RQ2. In contrast, the random flipping attack employs a budget of 15 but achieves $\Delta\text{GEA}$ of almost zero on all datasets, indicating it is ineffective. Meanwhile, the random rewiring attack uses a budget of 16 and achieves a very large $\Delta\text{GEA}$ (ranging from $0.293$ to $0.596$). However, this comes at the cost of large $\Delta\text{Label}$ (ranging from $0.159$ to $0.274$). In other words, the predicted labels of many nodes have been changed after the attacks, and this violates our objective.
5.3 Experiments to Answer RQ3: Attack Transferability
Answer to RQ3. From Table 2, we can see that some post-hoc explainers, including Grad, GuidedBP, PGMExplainer, and IG, suffer from poor explanation accuracy on the original graphs, even giving lower accuracy than the Random Explainer. As a result, investigating the attack transferability on these explainers may not be meaningful. On the contrary, like PGExplainer, explainers such as GradCAM, GNNExplainer, and SubgraphX (gray) can achieve good performance on most original graphs. For those, the attacks optimized for PGExplainer are also harmful, leading to a dramatic degradation of explanation accuracy. For instance, the explanation accuracy of GNNexplainer decreases from 0.711 to 0.397 on Syn2. This demonstrates the transferability of the attacks generated by GXAttack.
Other Analysis. We performed a sensitivity analysis for GXAttack in terms of maximal attack budget, maximal training epochs, and the loss trade-off hyperparameter. The results and analysis are deferred to Appendix I due to space reasons. Moreover, we also conducted some property analysis of GXAttack in terms of the prediction confidence, and node degree in Appendix H.
Table 2: Transferability of attacks on selected synthetic datasets with a typical run. The results are consistent on other datasets and thus are deferred to Table 12 to save space. We report the explanation accuracy (i.e., GEA) on the original graph $\rightarrow$ explanation accuracy on the perturbed graph.
Explainer Syn1 Syn2 Syn3 Syn4
PGExplainer 0.728 $\rightarrow$ 0.510 0.648 $\rightarrow$ 0.347 0.485 $\rightarrow$ 0.304 0.471 $\rightarrow$ 0.311
Grad 0.025 $\rightarrow$ 0.065 $0$rightarrow 0.001 $0$rightarrow 0.000 0.003 $\rightarrow$ 0.001
GradCAM 0.745 $\rightarrow$ 0.551 0.708 $\rightarrow$ 0.397 0.510 $\rightarrow$ 0.300 0.511 $\rightarrow$ 0.310
GuidedBP $0$rightarrow 0.000 $0$rightarrow 0.000 $0$rightarrow 0.000 $0$rightarrow 0.000
IG 0.118 $\rightarrow$ 0.104 0.134 $\rightarrow$ 0.103 0.136 $\rightarrow$ 0.101 0.109 $\rightarrow$ 0.092
GNNExplainer 0.738 $\rightarrow$ 0.550 0.711 $\rightarrow$ 0.397 0.539 $\rightarrow$ 0.341 0.514 $\rightarrow$ 0.327
SubgraphX 0.113 $\rightarrow$ 0.015 0.532 $\rightarrow$ 0.375 0.296 $\rightarrow$ 0.141 0.417 $\rightarrow$ 0.282
PGMExplainer 0.015 $\rightarrow$ 0.064 0.046 $\rightarrow$ 0.039 0.031 $\rightarrow$ 0.034 0.036 $\rightarrow$ 0.041
RandomExplainer 0.184 $\rightarrow$ 0.146 0.170 $\rightarrow$ 0.096 0.135 $\rightarrow$ 0.102 0.124 $\rightarrow$ 0.090
6 Related Work
We only briefly revisit closely related work due to space constraints. See Appendix F for more detail.
GNN Explanations. Yuan et al. (2022); Li et al. (2022); Kakkad et al. (2023) perform comprehensive surveys and propose excellent taxonomies of GNN explanation methods. Specifically, GNN explanations methods mainly include: 1) Gradients-based methods: Grad (Simonyan et al., 2013), GradCAM (Pope et al., 2019), GuidedBP (Baldassarre and Azizpour, 2019), Integrated Gradients (IG) (Sundararajan et al., 2017); 2) Decomposition-based methods: GraphLRP (Schwarzenberg et al., 2019), GNN-LRP (Schnake et al., 2021); 3) Surrogate-based methods: GraphLIME (Huang et al., 2022), PGM-Explainer (Vu and Thai, 2020); 4) Generation-based methods: XGNN (Yuan et al., 2020); 5) Perturbation-based methods: GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), GraphMask (Schlichtkrull et al., 2020), SubgraphX (Yuan et al., 2021); and 6) CF-based methods: CF-GNNExplainer (Lucic et al., 2022), RCExplainer (Bajaj et al., 2021).
Robustness/Stability of GNN Explanations. Yuan et al. (2022); Agarwal et al. (2022) point out that a reliable GNN explainer should exhibit stability, namely minor perturbations that do not impact the model predictions should not largely change the explanations. Particularly, Agarwal et al. (2022) theoretically analyze the stability of three GNN explainers, including vanilla Grad, GraphMask and GraphLIME, for which they derive upper bounds on the instability of explanations. Moreover, Bajaj et al. (2021); Fang et al. (2024); Wang et al. (2024); Fan et al. (2023) are closely related but are different from our work (as compared in Section 1 and detailed in Appendix F).
Adversarial Robustness of GNNs. Starting with the seminal works (Zügner et al., 2018; Dai et al., 2018), that both proposed adversarial attacks on GNNs, a rich literature formed in the realm of GNN-specific adversarial attacks, defenses, and certifications (Jin et al., 2021; Günnemann, 2022). Most attacks primarily focus on altering predictions via perturbations of the discrete graph structure (edge insertions/deletions), either between existing nodes (Zügner et al., 2018) or via the insertion of new (adversarial) nodes (Zou et al., 2021). Optimization-based attacks either operate globally (number of edge insertions/deletions) (Xu et al., 2019; Geisler et al., 2021), alter a local neighborhood (Zügner et al., 2018), or can handle a combination of both constraints (Gosch et al., 2023). While adversarial attacks on GNN explanation methods may benefit from optimization methods over the discrete graph structure, a careful discussion and derivation of appropriate attack objectives and metrics was missing.
7 Conclusions
GNN explanations for enhancing transparency and trust of graph neural networks are becoming increasingly important in decision-critical domains. Our research reveals that existing GNN explanation methods are vulnerable to adversarial perturbations that strongly change the explanations without affecting the predictions. This vulnerability undermines the reliability of these methods in crucial settings. Specifically, we introduced GXAttack, an optimization-based adversarial method to assess the robustness of post-hoc GNN explanations. GXAttack’s effectiveness underscores the need for new evaluation standards that incorporate adversarial robustness as a fundamental aspect. Future research should focus on developing explanation methods that are both interpretable and robust against adversarial attacks, as enhancing the stability of GNN explanations is crucial for ensuring their consistency and reliability in practical applications.
The limitations and broader impact of our work are discussed in Appendices J and  K, respectively.
Acknowledgements
Zhong Li and Matthijs van Leeuwen: this publication is part of the project Digital Twin with project number P18-03 of the research programme TTW Perspective, which is (partly) financed by the Dutch Research Council (NWO). Simon Geisler: This research was supported by the Helmholtz Association under the joint research school “Munich School for Data Science – MUDS".
Appendix for paper submission Explainable Graph Neural Networks Under Fire
Appendix A Revisiting PGExplainer
For completeness, we revisit PGExplainer as follows. For brevity, we denote $\mathcal{G}(v)$ as $\mathcal{G}$ (and $\mathcal{G}_{s}(v)$ as $\mathcal{G}_{s}$ analogously) without causing confusions throughout this subsection.
Formalisation. Given a graph $\mathcal{G}$ and the target node $v$, to explain why $f_{\theta}(\mathcal{G}(v))=c$, PGExplainer will provide a local explanation (i.e. explanatory subgraph) $\mathcal{G}_{s}(v)=(\mathbf{A}_{s},\mathbf{X}_{s})$, where $\mathbf{A}_{s}$ is the subgraph structure and $\mathbf{X}_{s}$ is the node features associated with nodes residing in this subgraph. Specifically, PGExplainer maximises the mutual information between the GNN classifier’s prediction $\mathbf{y}$ and the explanatory subgraph $\mathcal{G}_{s}$:
$\underset{\mathcal{G}_{s}}{\max}~{}\text{MI}(\mathbf{y},\mathcal{G}_{s}):=H(\mathbf{y})-H(\mathbf{y}|\mathcal{G}_{s})$ (17)
where $H(\mathbf{y})$ is the entropy of $f_{\theta}(\mathcal{G})$, and $H(\mathbf{y}|\mathcal{G}_{s})$ is the conditional entropy of the GNN’s prediction on the explanatory graph, namely $f_{\theta}(\mathcal{G}_{s})$. As $\theta$ and $\mathcal{G}$ are fixed at the explanation stage (i.e. $\mathbf{y}$ is also fixed), maximising Eq 17 is equivalent to minimising $H(\mathbf{y}|\mathcal{G}_{s})$.
Relaxations. However, directly minimising $H(\mathbf{y}|\mathcal{G}_{s})$ is intractable as there are $\mathcal{O}(2^{|\mathcal{E}|})$ candidates for $\mathcal{G}_{s}$. Therefore, they considered three relaxations: 1) they assume that the explanatory subgraph $\mathcal{G}_{s}$ is a Gilbert random graph, namely the selections of edges from $\mathcal{G}$ are conditionally independent to each other. As a result, the probability of graph $\mathcal{G}$ can be factorised as $P(\mathcal{G})=\prod_{(i,j)\in\mathcal{E}}P(e_{ij})$. Moreover, they instantiate $P(e_{ij})$ using a Bernoulli distribution $B(\lambda_{ij})$, namely $e_{ij}\sim B(\lambda_{ij})$ with $P(e_{ij}=1)=\lambda_{ij}$. As a result, Eq 17 is equivalent to:
$\underset{\mathcal{G}_{s}}{\min}~{}H(\mathbf{y}|\mathcal{G}_{s})=\underset{\mathcal{G}_{s}}{\min}~{}\mathbb{E}_{\mathcal{G}_{s}}[H(\mathbf{y}|\mathcal{G}_{s})]\approx\underset{\Lambda}{\min}~{}\mathbb{E}_{\mathcal{G}_{s}\sim q(\Lambda)}[H(\mathbf{y}|\mathcal{G}_{s})],$ (18)
where $q(\Lambda)$ describes the distribution of explanatory subgraph $\mathcal{G}_{s}$ that is parameterized by $\lambda$’s in $\Lambda$. 2) they relax the edge weights from binary variables $\{0,1\}$ to continuous variable $[0,1]$ using reparameterization trick. As a result, they can utilise gradient-based methods to optimize the objective function. More specifically, the continuous edge weight $\hat{e}_{ij}$ is calculated as $\hat{e}_{ij}=sigmoid[(\ln(\epsilon)-\ln(1-\epsilon)+\omega_{ij})/\tau]$, where $\epsilon\sim\text{U}(0,1)$, $\tau$ is the temperature hyperparameter, and $\omega_{ij}$ is the learnable parameter in $\Omega$. Hence, Eq 18 is equivalent to:
$\underset{\Omega}{\min}~{}\mathbb{E}_{\epsilon\sim\text{U}(0,1)}[H(\mathbf{y}|\mathcal{G}_{s})].$ (19)
3) for efficiency, they replace conditional entropy $H(\mathbf{y}|\mathcal{G}_{s})$ with cross entropy $H(\mathbf{y},\mathbf{y}_{s})$, and further apply Monte Carlo to approximate it:
$\underset{\Omega}{\min}~{}\frac{-1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\mathcal{G}_{s}^{(k)})),$ (20)
where $K$ is the number of sampled explanatory subgraph $\mathcal{G}_{s}^{(k)}$, and $\theta$ is the set of parameters in trained $f_{\theta}(\cdot)$.
Generalisations. If we consider $\Omega$ in Eq 20 as independent variables, then we have to retrain it each time when we want to explain an unseen instance. To mitigate this issue, they employ a parameterized network to learn $\Omega$, namely $\Omega=h_{\Phi}(\mathcal{G},\mathbf{Z})$, where $\mathbf{Z}$ is the learned node embedding matrix before inputting to the downstream tasks (namely the result of component 1 in GNN classifier). Besides, $h(\cdot)$ is the explanation network that is shared by all edges, and $\Phi$ is the set of its parameters. As a result, in the collective setting Eq 20 becomes:
$\underset{\Phi}{\min}~{}\sum_{i\in\mathcal{I}}\sum_{k=1}^{K}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G}^{(i)})\ln(P_{\theta}(y=c|\mathcal{G}_{s}^{(i,k)})),$ (21)
where $\mathcal{G}^{(i)}$ is the $i$-th input graph from a set $\mathcal{I}$, and $\mathcal{G}_{s}^{(i,k)}$ is its $k$-th sampled explanatory subgraph.
Finally, they implement the explanatory network $h(\cdot)$ as a MLP, which returns the edge importance as explanations. For node classification, the explanation is $\omega_{ij}=\text{MLP}_{\Phi}([\mathbf{z}_{i},\mathbf{z}_{j};\mathbf{z}_{v}])$ for node $v$.
Appendix B Projected Gradient Descent
After each gradient update $\tilde{\mathbf{P}}_{t}\in\mathbb{R}^{n\times n}$ although it must be in $[0,1]^{n\times n}$ for its interpretation as Bernoulli random variable. Moreover, we must ensure that sampling $\mathbf{P}_{t}\in\{0,1\}^{n\times n}\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{t})$ will yield a discrete perturbation $\hat{\mathbf{A}}=\mathbf{A}\oplus\mathbf{P}_{t}$ obeying the admissible perturbations $\|\mathbf{A}-\hat{\mathbf{A}}\|_{0}<B$ with sufficient likelihood. Following Xu et al. (2019); Geisler et al. (2021), we ensure that $\mathbb{E}_{\mathbf{P}\sim\text{Bernoulli}(\tilde{\mathbf{P}_{t}})}[\|\mathbf{P}\|_{0}]\leq B$ and then obtain the final perturbation form a small set of samples. To guarantee that $\tilde{\mathbf{P}}_{t}$ parametrizes a valid distribution and that we obey the budget in expectation, we employ the projection $\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})\leq B]}(\tilde{\mathbf{P}}_{t})$.
The space spanned by $\tilde{\mathbf{P}}_{t}\in[0,1]^{n\times n}$ describes a $n\times n$ dimensional hypercube intersected with the region below the $B$-simplex defining $\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})=B]$. The region below the $B$-simplex can be expressed as $\vec{1}^{\top}\tilde{\mathbf{P}}\vec{1}=\sum_{i,j}\tilde{\mathbf{P}}_{i,j}\leq B$. Thus, after each gradient update, we solve
$\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}})\leq B]}(\tilde{\mathbf{P}})=\arg\min_{\tilde{\mathbf{S}}}\|\tilde{\mathbf{P}}-\tilde{\mathbf{S}}\|_{F}^{2}\qquad\text{subject to }\;$ $\sum_{i,j}\tilde{S}_{i,j}\leq B$ (22)
$\;\tilde{\mathbf{S}}\in[0,1]^{n\times n}$
for clipping to the hyperplane and projecting back towards the simplex using a bisection search. See Xu et al. (2019) for details and the derivation.
Appendix C Complexity Analysis of the GXAttack Algorithm
Algorithmic Components
  1. 1.
    Initial Predictions and Explanations:
    • $f_{\theta}(\mathbf{A},\mathbf{X})$: Predicting using the GNN on the graph $\mathcal{G}$. The complexity depends on the GNN architecture but typically involves matrix multiplications between the adjacency matrix $\mathbf{A}$ and the feature matrix $\mathbf{X}$.
    • $g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X}))$: Generating explanations based on the GNN predictions. The complexity is generally proportional to that of the GNN prediction if the explainer works by perturbing inputs and observing changes in output.
  2. 2.
    Loop for Gradient Updates:
    • Prediction and Explanation on Perturbed Graph: This operation is similar in complexity to the initial prediction and explanation but is repeated $T$ times (number of epochs).
    • Gradient Computation and Update:
      • Calculating the gradient $\nabla_{\tilde{\mathbf{P}}_{t-1}}\mathcal{L}$ involves computing derivatives of the loss function with respect to the perturbation matrix $\tilde{\mathbf{P}}$. This can be expensive as it depends on the sensitivity of both the GNN predictions and explanations to changes in the adjacency matrix.
      • Updating $\tilde{\mathbf{P}}_{t}$ involves a simple addition and a projection operation. The projection step is constrained by the perturbation budget $B$, but the actual operation is typically linear with respect to the number of potential edges $n^{2}$.
  3. 3.
    Binary Perturbation Sampling:
    • Sampling from a Bernoulli distribution for each potential edge (pair of nodes), which has a complexity proportional to $n^{2}$.
Computational Complexity
We discuss the dense case here and refer to the summary below for a discussion of GNNs/explainers implemented with sparse matrices.
  1. 1.
    Initial Operations: $O(f_{\theta}+g_{\lambda})$, typically $O(n^{2})$ for a dense GNN and explainer (sometimes even $O(n^{3})$).
  2. 2.
    Each Epoch:
    • Prediction and explanation on perturbed graph: $O(2T\cdot(f_{\theta}+g_{\lambda}))$.
    • Gradient computation and update: Dependent on the efficiency of backpropagation through the GNN, but generally $O(T\cdot n^{2})$ for gradient computation and a similar order for updates.
  3. 3.
    Sampling Step: $O(n^{2})$.
Space Complexity
  1. 1.
    Space for Storing Matrices: Storing $\mathbf{A}$, $\mathbf{X}$, $\tilde{\mathbf{P}}$ each requires $O(n^{2})$ space.
  2. 2.
    Additional Space for Operations: Space required for intermediate gradients and outputs during predictions, typically $O(n^{2})$.
Summary
Since the explainer uses dense matrices, we also choose an attack operating on all $n^{2}$ edges. Thus, the total complexity $O((T+1)\cdot(f_{\theta}+g_{\lambda})+n^{2})$ since we have $T$ attack steps plus the initial predictions as well as explanations and, last, we sample the final perturbations. Assuming that $O(f_{\theta})=O(g_{\lambda})=O(n^{2}d+nd^{2})$ and $d\ll n$, this yields a total time complexity of $O(n^{2})$ with number of attack iterations $T\ll n$, and number of nodes $n$. This is the same asymptotic complexity as the dense GNN $f_{\theta}$ and explainer $g_{\lambda}$.
Assuming GNN and explainer are in $O(f_{\theta})=O(g_{\lambda})=O(Ed+nd^{2})$ with number of edges $E$, a dense attack would dominate the overall complexity ($O(n^{2})$). An immediate remedy would yield PRBCD (Geisler et al., 2021) since its complexity is $O(B+E)$ with edge flipping budget $B$ and all additional components (e.g., loss function) could be implemented in $O(E)$ with sparse matrices. This yields an overall time and space complexity of $O(E)$, assuming $B\ll E$ and hidden dimensions $d\ll n$, which is the same asymptotic complexity as the sparse GNN $f_{\theta}$ and explainer $g_{\lambda}$.
Appendix D Datasets Description
Overall, we consider the following synthetic datasets, which are summarised in Table 3: 1) synthetic datasets with different size via varying the following hyper-parameters in ShapeGGen: num_subgraphs for the number of subgraphs and subgraph_size for the number of expected nodes in each subgraph; 2) synthetic datasets with different types of explanatory graphs: shape as “circle" or “house"; and 3) synthetic datasets with different node degrees: prob_connection with different values.
Table 3: Summary of synthetic datasets with ground-truth explanations. #Nodes and #Edges represent the number of expected nodes and edges, respectively. We do not consider larger datasets for two reasons: 1) explanation accuracy will be very low ($<0.3$) for most explainers, and it is not meaningful to attack low accuracy explainers; 2) computation will be slow for the attacker (and the explainer).
Dataset Syn1 Syn2 Syn3 Syn4 Syn5 Syn6 Syn7
Shape House House House Circle House House House
#Subgraphs 10 50 50 50 50 50 100
Subgraph Size 6 6 10 10 10 10 10
P(Connection) 0.3 0.06 0.06 0.06 0.12 0.20 0.06
#Classes 2 2 2 2 2 2 2
#Nodes 59 299 521 511 489 492 1018
#Edges 164 970 1432 1314 1664 1898 3374
Average Degree 2.8 3.2 2.7 2.6 3.4 3.9 3.3
Max Budget 10 15 15 15 15 15 15
Appendix E Presumptions Checking Results
Table 4: Complete results on Syn1. O. GEA means explanation accuracy on original graph and P. GEA indicates explanation accuracy on perturbed graph. Moreover, $\Delta$GEA is the explanation accuracy change after perturbation, Simcos is the cosine similarity between explanations (before and after perturbations), and #Pert. represents the number of flipped edges. Besides, $\Delta$Label is the average change of predicted labels, while $\Delta$Prob denotes the average absolute change of predicted probability (of the original predicted class). “G5" is the set of nodes of which the original prediction confidence located in [0.5,0.6). Similarly, “G6" for [0.6,0.7), “G7" for [0.7,0.8), “G8" for [0.8,0.9), “G9" for [0.9,1.0], and “All" for all nodes.
Confidence Group G5 G6 G7 G8 G9 All
O. GEA 0.696 0.763 0.726 0.775 0.683 0.728
P. GEA ()($\downarrow$)( ↓ ) 0.403 0.463 0.485 0.584 0.679 0.560
$\Delta$GEA ()($\uparrow$)( ↑ ) 0.293 0.299 0.241 0.190 0.004 0.168
Simcos ()($\downarrow$)( ↓ ) 0.817 0.868 0.898 0.937 0.987 0.923
#Pert ()($\downarrow$)( ↓ ) 6.5 5.0 4.5 2.7 0.2 3.0
$\Delta$Label ()($\downarrow$)( ↓ ) 0 0 0 0 0 0
$\Delta$Prob ()($\downarrow$)( ↓ ) 0.327 0.204 0.147 0.065 0.002 0.108
Table 5: Complete results on Syn2. See Table 4 for the meaning of row names and column names.
Confidence Group G5 G6 G7 G8 G9 All
O. GEA 0.694 0.739 0.690 0.694 0.546 0.648
P. GEA ()($\downarrow$)( ↓ ) 0.329 0.396 0.390 0.387 0.381 0.347
$\Delta$GEA()($\uparrow$)( ↑ ) 0.365 0.293 0.300 0.307 0.165 0.269
Simcos()($\downarrow$)( ↓ ) 0.958 0.960 0.971 0.979 0.991 0.977
#Pert.()($\downarrow$)( ↓ ) 7.1 6.6 6.2 6.3 3.8 5.6
$\Delta$Label()($\downarrow$)( ↓ ) 0 0 0 0 0 0
$\Delta$Prob()($\downarrow$)( ↓ ) 0.353 0.268 0.181 0.103 0.003 0.143
(a) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn1.
(b) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn2.
(c) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn3.
(d) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn4.
(e) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn5.
(f) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn6.
(g) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn7.
Figure 2: Comparative analysis of explanation accuracy with respect to prediction correctness on all synthetic datasets. The results show that explanation accuracy is not necessarily correlated with the prediction correctness. Particularly, we did not observe that wrongly predicted nodes tend to have lower explanation accuracy.
Table 6: Complete results on Syn3. See Table 4 for the meaning of row names and column names.
Confidence Group G5 G6 G7 G8 G9 All
O. GEA 0.484 0.496 0.479 0.510 0.440 0.480
P. GEA()($\downarrow$)( ↓ ) 0.305 0.303 0.279 0.335 0.310 0.306
$\Delta$GEA()($\uparrow$)( ↑ ) 0.179 0.193 0.200 0.175 0.130 0.179
Simcos()($\downarrow$)( ↓ ) 0.966 0.973 0.962 0.981 0.989 0.978
#Pert.()($\downarrow$)( ↓ ) 8.0 7.4 6.6 5.8 4.7 6.7
$\Delta$Label()($\downarrow$)( ↓ ) 0.008 0 0 0 0 0.002
$\Delta$Prob()($\downarrow$)( ↓ ) 0.242 0.206 0.166 0.099 0.039 0.165
Table 7: Complete results on Syn4. See Table 4 for the meaning of row names and column names.
Confidence Group G5 G6 G7 G8 G9 All
O. GEA()($\downarrow$)( ↓ ) 0.500 0.500 0.493 0.495 0.416 0.471
P. GEA()($\downarrow$)( ↓ ) 0.309 0.314 0.319 0.326 0.296 0.311
$\Delta$GEA()($\uparrow$)( ↑ ) 0.192 0.187 0.179 0.168 0.120 0.159
Simcos()($\downarrow$)( ↓ ) 0.973 0.972 0.962 0.979 0.989 0.980
#Pert.()($\downarrow$)( ↓ ) 6.9 6.8 6.4 6.2 3.9 5.7
$\Delta$Label()($\downarrow$)( ↓ ) 0 0 0 0 0 0
$\Delta$Prob()($\downarrow$)( ↓ ) 0.333 0.267 0.177 0.111 0.030 0.149
Table 8: Complete results on Syn5. See Table 4 for the meaning of row names and column names.
Confidence Group G5 G6 G7 G8 G9 All
O. GEA 0.456 0.470 0.443 0.461 0.430 0.450
P. GEA()($\downarrow$)( ↓ ) 0.252 0.264 0.282 0.305 0.336 0.293
$\Delta$GEA()($\uparrow$)( ↑ ) 0.204 0.207 0.161 0.157 0.094 0.157
Simcos()($\downarrow$)( ↓ ) 0.967 0.971 0.975 0.979 0.984 0.976
#Pert.()($\downarrow$)( ↓ ) 7.5 7.6 7.0 6.6 5.2 6.6
$\Delta$Label()($\downarrow$)( ↓ ) 0.013 0 0 0 0 0.002
$\Delta$Prob()($\downarrow$)( ↓ ) 0.287 0.226 0.147 0.085 0.031 0.138
Table 9: Complete results on Syn6. See Table 4 for the meaning of row names and column names.
Confidence Group G5 G6 G7 G8 G9 All
O. GEA 0.415 0.412 0.407 0.382 0.291 0.332
P. GEA()($\downarrow$)( ↓ ) 0.260 0.249 0.239 0.259 0.217 0.231
$\Delta$GEA()($\uparrow$)( ↑ ) 0.155 0.162 0.168 0.123 0.074 0.101
Simcos()($\downarrow$)( ↓ ) 0.975 0.975 0.979 0.986 0.984 0.989
#Pert.()($\downarrow$)( ↓ ) 7.3 7.8 7.5 7.0 5.0 5.9
$\Delta$Label()($\downarrow$)( ↓ ) 0 0 0 0 0 0
$\Delta$Prob()($\downarrow$)( ↓ ) 0.348 0.270 0.189 0.106 0.029 0.092
Table 10: Complete results on Syn7. See Table 4 for the meaning of row names and column names.
Confidence Group G5 G6 G7 G8 G9 All
O. GEA 0.428 0.415 0.406 0.405 0.401 0.409
P. GEA()($\downarrow$)( ↓ ) 0.266 0.250 0.294 0.263 0.278 0.263
$\Delta$GEA ()($\uparrow$)( ↑ ) 0.161 0.165 0.156 0.142 0.123 0.146
Simcos()($\downarrow$)( ↓ ) 0.986 0.983 0.989 0.991 0.992 0.990
#Pert.()($\downarrow$)( ↓ ) 8.2 7.9 7.5 7.1 6.1 7.2
$\Delta$Label()($\downarrow$)( ↓ ) 0 0 0 0 0 0
$\Delta$Prob()($\downarrow$)( ↓ ) 0.319 0.253 0.169 0.097 0.035 0.147
(a) Explanation accuracy changes after attacks for nodes with prediction confidence ($0.60.5\leq p<$ 0.6). One can see that the attacks tend to be more successful if the original explanation accuracy is higher.
(b) Explanation accuracy changes after attacks for nodes with different prediction confidence.
Figure 3: Comparison of explanation accuracy changes after attacks across different prediction confidences on Syn2. The starting point of each directed line segment represents the original accuracy, the end point (i.e., the arrow position) represents the accuracy after perturbation, and the length of the line segment represents the change in explanation accuracy.
Appendix F Related Work in More Detail
Related work includes adversarial attacks and defenses on graphs, GNN explanations, attacks and defenses of traditional XAI methods, and stability of GNN explanation methods.
F.1 Adversarial Attacks and Defenses on Graphs
Graph adversarial attacks attempt to deliberately manipulate or deceive graph learning models by making small, often imperceptible/unnoticeable perturbations to the input graph’s structure (i.e., edge/node additions/deletions) and/or attributes (i.e., attribute perturbations). Particularly, most existing adversarial attacks and defenses methods focus on node-level operations and tasks Jin et al. (2021). Furthermore, these methods can be roughly categorised into: 1) white-box attacks, 2) gray-box attacks, and 3) black-box attacks according to the attacker’s knowledge.
Specifically, white-box attacks usually assume that attackers can calculate gradient via model parameters, which are generally not feasible in real-world applications. Meanwhile, gray box attacks usually assume that they have access to the training data and corresponding labels (node-level or graph-level). On this basis they usually train a surrogate model to mimic the real model (not accessible), and then attack the surrogate model. Black box attacks can only access the adjacency matrix, node attribute matrix, and the output of the victim GNN model (model parameters, ground-truth labels and predicted probabilities are not available). Most black-box attacks are constructed based on the attack transferability assumption (namely attacks are effective cross different models) and perform queries from victim model.
F.2 GNN Explanations
Yuan et al. (2022); Li et al. (2022); Kakkad et al. (2023) perform comprehensive surveys and propose excellent taxonomies of GNN explanation methods. GNN explanations methods can be divided into the following categories based on Yuan et al. (2022):
  • Gradients-based methods: Vanilla Grad (Simonyan et al., 2013), GradCAM (Pope et al., 2019), GuidedBP (Baldassarre and Azizpour, 2019), Integrated Gradients (IG) (Sundararajan et al., 2017);
  • Decomposition-based methods: GraphLRP (Schwarzenberg et al., 2019), Excitation BP, GNN-LRP (Schnake et al., 2021);
  • Surrogate-based methods: GraphLIME (Huang et al., 2022), PGM-Explainer (Vu and Thai, 2020);
  • Generation-based methods: XGNN (Yuan et al., 2020);
  • Perturbation-based methods: GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), GraphMask (Schlichtkrull et al., 2020), SubgraphX (Yuan et al., 2021);
  • CF-based methods: CF-GNNExplainer (Lucic et al., 2022), RCExplainer (Bajaj et al., 2021).
Particularly, they are summarised in Table 11.
Table 11: Summary of GNN explanation methods.
Method Venue Datasets Metrics Code
GNNExplainer NIPS’19 Synthetic (BA-Shapes, BA-Community, Tree-Cycles, Tree-Grid) for node classification; Mutag and Reddit-Binary for graph classification Explanation Accuracy Github
GraphLIME TKDE’22 Cora, PubMed for node classification Noisy features ratio Github
XGNN KDD’20 Synthetic (Is_Acyclic) and Mutag for graph classification Visualisation Github
PGExplainer NIPS’20 Synthetic (BA-Shapes, BA-Community, Tree-Cycles, Tree-Grid) for node classification; Synthetic (BA-2motifs) and Mutag for graph classification Explanation Accuracy; Visualisation Github
PGM-Explainer NIPS’20 Synthetic, Bitcoin-Alpha and Bitcoin-OTC for node classification; SuperPixel-Graph for graph classification Explanation Accuracy; Human-subjective test Github
Gradients CVPR’19 Visual Genome scene-graphs, BBBP, BACE, and TOX21 for graph classification Fidelity, Contrastivity, Sparsity Github
SubgraphX ICML’21 Synthetic (BA-Shapes) for node classification; Synthetic (BA-2Motifs), Mutag, BBBP, Graph-SST2 for graph classification Fidelity, Sparsity, Visualisation Github
CF-GNNExplainer AIST’22 Synthetic (BA-Shapes, Tree-Cycles, Tree-Grid) for node classification; Fidelity, Sparsity, Explanation Size, Accuracy Github
RCExplainer NIPS’21 Synthetic (BA-Shapes, BA-Community, Tree-Cycles, Tree-Grid) for node classification; Synthetic (BA-2motifs), Mutag, NCI1 for graph classification; Fidelity, Robustness, Accuracy, Efficiency Github
V-InFoR NIPS’23 Synthetic (BA-3motifs), Mutag, Ogbg-molhiv, Ogbg-ppa for graph classification; Sufficient Accuracy, Necessary Accuracy, $F_{NS}$ Score Github
GraphSVX ECML’21 Synthetic (BA-Shapes, BA-Community, Tree-Cycles, Tree-Grid) for node classification; Synthetic (BA-2motifs) and Mutag for graph classification; Accuracy, Robustness, Certainty Github
F.3 Attacks and Defenses of Traditional XAI Methods
It has been shown that traditional XAI methods (namely those designed to explain deep neural networks for image and text data) are susceptible to malicious perturbations (Heo et al., 2019; Zhang et al., 2020; Ghorbani et al., 2019). Please refer to Baniecki and Biecek (2024) for a comprehensive survey.
F.4 Robustness of GNN Explanations
OAR. Fang et al. (2024) present a novel metric for evaluating GNN explanations, termed Out-Of-Distribution-resistant Adversarial Robustness (OAR). OAR aims to solve the limitations of current removal- and generative-based evaluations. More concretely, they evaluate post-hoc explanation subgraphs by computing their robustness under attack, which consists of three steps: 1) they formulate the adversarial robustness tailored for GNN explanations problem, which is the minimum adversarial perturbation on the structure of complementary subgraph (i.e., removing explanatory subgraph from the original graph); In other words, a good explanation subgraph will yield high “such-defined” adversarial robustness; 2) they introduce a tractable and easy-to-implement objective of adversarial robustness for GNN explanations; after relaxation, the objective becomes a constrained graph generation problem (namely random perturbations) rather than an adversarial attack problem; 3) they present an OOD reweighting block (which is instantiated using a graph anomaly detector VGAE) that aims to confine the evaluation on original data distribution. The OOD score is used as the weight of each prediction when computing the expectation of generated graph’s prediction.
They evaluate the performance of post-hoc explainers including GNNExplainer, PGExplainer, Refine, SA, CXPlain, and GradCAM, on BA3, TR3, Mutag and MNIST datasets for graph classification task. They consider ground-truth explanations, consistency with metric that is defined using ground-truth explanations, and consistency with human intuition. This work is an initial attempt to explore evaluation metrics for GNN explanations from the perspectives of adversarial robustness and resistance to OOD.
V-InFoR. (Wang et al., 2024) point out that “existing GNN explainers are not robust to the structurally corrupted graphs, namely graphs with noisy or adversarial edges." This study for the first time investigates the negative effect of minor structural corruptions (which do not change the predictions) and severe structural corruptions (which change the predictions) on GNN explainers. Particularly, they provide quantitative evidence that existing GNN explainers are fragile to structurally corrupted graphs. They evaluate 6 GNN explainers (GradCAM, IG, PGM-Explainer, GNNExplainer, PGEXplainer, ReFine) on BA-3Motifs dataset under two settings: 1) minor corruptions where they select the perturbed graphs whose predictions are not changed after randomly flipping 20% edges; 2) severe corruptions where they employ adversarial attack algorithm GRABNEL to perturb at most 10% edges to change the predictions. Their results shows that minor corruptions lead to an average of 31.15% explanation performance degradation, and severe corruptions lead to 40.64%.
Particularly, V-InfoR consists of two modules: 1) a robust graph representation learner that aims to learn a variational distribution that contains common information shared a large amount of graphs; 2) an adaptive explanation generator that samples explanatory subgraph from the variational distribution.
RCExplainer. Bajaj et al. (2021) aim to find a small set of edges of the input graph such that the prediction result will substantially change if we remove those edges. Specifically, they propose RCExplainer to generate robust counter-factual explanations for GNNs in two steps: 1) they employ a set of decision regions to model the decision logic of a GNN, where each decision region (corresponding to a class) determines the predictions on multiple graphs that are predicted to be the same class; 2) they leverage a DNN model to explore the decision logic (namely the linear decision boundaries of the decision regions), and thereby extract robust counter-factual explanations as a small set of edges of the input graph. Particularly, the linear decision boundaries of the decision region (of a specific class) capture the common decision logic on all graphs inside the decision region, making it not easy to overfit the noise of some individual graphs. This method is only applicable to GNNs that belong to Piecewise Linear Neural Networks (e.g., ReLU, Leaky ReLU, and Maxout).
They point out (for the first time) that most methods such as PGExplainer, Gradients, Graph Attentions, and GNNExplainer are not robust to noise, namely slightly random perturbations on input graph may cause that the generated explanations substantially change while the predicted labels keep unchanged. (They argue that this is because the explanation of every individual prediction is independently optimized to maximize its correlation with the prediction, and thus can easily cause overfitting to noise in the data.) Moreover, they indicate that these explanation methods are usually not counter-factual: “removing the explanation subgraph from the input graph does not necessarily change the prediction result due to the non-convexity of GNNs." However, they did not systematically investigate this phenomena.
GEAttack. Fan et al. (2023) empirically demonstrate that GNN explanation methods can be employed as tools to detect the adversarial perturbations on graphs. On this basis, they propose an attack framework dubbed GEAttack that can jointly attack both GNN and its explanations. Specifically, they formulate GEAttack as a bi-level optimization problem: 1) they mimic the GNN explanation method optimization process to obtain graph explanations in the inner loop; 2) they compute the gradient of the attack objective w.r.t. the explanations in the outer loop. Moreover, they considered two GNN explanation methods, including GNNExplainer and PGExplainer.
Appendix G Training Protocols and Compute Resources
G.1 GNN Predictor Training Protocols
To avoid pitfalls caused by weak GNN predictor Faber et al. (2021), we should train the GNN predictor close to its maximum possible performance. Moreover, when training the GNN predictor, we employ an Adam optimizer where the learning rate is 1e-2, weight decay is 1e-5, the number of epochs is 300, and the hidden dimension is 32.
G.2 GNN Explainer Training Protocols
We followed the authors’ recommendations for setting the hyperparameters of GNN explanation methods. We select top-$k~{}(k=25\%)$ important edges (also top-$k$ nodes and node features if necessary) to generate explanations for all GNN explainers. Particularly, when training the PGExplainer, the number of training epochs is 10.
G.3 GNN Explanation Attacker Training Protocols
We set the trade-off hyperparameter in the final loss as $0.1$ such that the magnitudes of the two loss terms are approximately the same. Besides, the learning rate in gradient ascent is set to $0.1$. Importantly, we set the maximal training epochs as 100 due to computation time reason, while further increasing this number can largely improve the attack performance (in terms of $\Delta$GEA). Moreover, we set the perturbation budget to 15 at maximum for synthetic data (note that we attempt to employ a hard minimum budget of 5 with 2000 maximum trials of Bernoulli sampling). We perform sensitivity analysis regarding these hyperparameters, and the results and analysis are detailed in Appendix I.
G.4 Explanation Accuracy Computation
The $get\_acc(\cdot)$ function of GraphXAI considers all non-zero edges as predicted positives (namely with binary code ‘1’) by default when computing JAC. Following Agarwal et al. (2023), we set the top 25% (of edges with the highest edge importance scores) as 1 and the rest as zeros in our evaluations when computing both GEA and cosine similarity.
G.5 Software and Hardware
The evaluated GNN explainers are implemented mainly based on GraphXAI  [^1] https://github.com/mims-harvard/GraphXAI (with MIT License)(Agarwal et al., 2023). Algorithms are implemented in Python 3.8 (using PyTorch Paszke et al. (2019) and PyTorch Geometric Fey and Lenssen (2019) libraries when applicable) and ran on internal clusters with AMD 128 EPYC7702 CPUs (with 512GB RAM). All experiments can be finished within 7 days if using 10 such machine instances. All code and datasets will be available on GitHub [^2] https://github.com/ZhongLIFR/GXAttack.
Appendix H Additional Results
H.1 Property Analysis Regarding Prediction Confidence
(a) Prediction confidence VS original explanation accuracy on Syn2.
(b) Prediction confidence VS attack performance (i.e. $\Delta\text{GEA}$) on Syn2.
(c) Prediction confidence VS probability change after attacks on Syn2.
Figure 4: Property analysis regarding prediction confidence on Syn2. Other datasets show consistent results and thus are omitted. From Figure 4(a), we can clearly see that nodes in ‘G9’ tend to have lower original explanation accuracy. As a result, as shown in Figure 4(b), the attack performance of GXAttack on these nodes (in ‘G9’) tend to be less successful (i.e., smaller $\Delta\text{GEA}$). This is because the nodes in ‘G9’ leave no much room for perturbations under the constraint that the predictions remain unchanged, namely much smaller probability changes as shown in Figure 4(c).
H.2 Property Analysis Regarding Node Degree
(a) Node degree VS original explanation accuracy on Syn2.
(b) Node degree VS attack performance (i.e. $\Delta\text{GEA}$) on Syn2.
(c) Node degree VS probability change after attacks on Syn2.
Figure 5: Property analysis regarding node degree on Syn2. Other datasets show consistent results and thus are omitted. From Figure 5(a), we can see that nodes with higher degree tend to have lower original explanation accuracy. Figure 5(b) shows that the attack performance of GXAttack on these nodes with higher degree tend to be less successful (i.e., smaller $\Delta\text{GEA}$). Possible reasons are that 1) the original explanation accuracy tends to be lower; and 2) small perturbations on nodes with high degree have limited impact on the predictions, namely resulting small probability changes as shown in Figure 5(c).
H.3 Attack Transferability Results
Table 12: Transferability of attacks on synthetic datasets. We report the explanation accuracy on original graph $\rightarrow$ explanation accuracy on perturbed graph. The results on Syn7 are omitted due to the fact that runtime exceeded a 10-day limit for a single trial.
Explainer Syn5 Syn6
PGExplainer 0.450 $\rightarrow$ 0.285 0.332 $\rightarrow$ 0.225
Grad $0$rightarrow 0.000 $0$rightarrow 0.000
GradCAM 0.455 $\rightarrow$ 0.278 0.414 $\rightarrow$ 0.241
GuidedBP $0$rightarrow 0.000 $0$rightarrow 0.000
IG 0.116 $\rightarrow$ 0.085 0.078 $\rightarrow$ 0.065
GNNExplainer 0.483 $\rightarrow$ 0.283 0.431 $\rightarrow$ 0.205
SubgraphX 0.394 $\rightarrow$ 0.312 0.408 $\rightarrow$ 0.305
PGMExplainer 0.038 $\rightarrow$ 0.043 0.037 $\rightarrow$ 0.040
RandomExplainer 0.128 $\rightarrow$ 0.100 0.133 $\rightarrow$ 0.086
Appendix I Sensitivity Analysis
We provide the results of sensitivity analysis in this section. Specifically, we perform analysis in terms of $\Delta$GEA, $\Delta$Prob and used budget when varying the following hyperparameters, respectively: 1) the loss trade-off hyperparameter; 2) the maximally allowed perturbation budget (Max Budget) in Figure 6; and 3) the maximal training epochs of attacker in Figure 7;
Trade-off Hyperparameter vs Attack Performance. Our sensitivity analysis shows that the attack performance (in term of $\Delta$GEA), $\Delta$Prob, and the used budget are nearly constant when we vary the trade-off hyper-parameter $\beta$’s value from $0.000001$ to 10. We omit the figures here.
Attack Budget vs Attack Performance.
Figure 6: Sensitivity analysis w.r.t. maximally allowed perturbation budget (Max Budget) on Syn2 using GXAttack. We can see that as the Max Budget increases from 1 to 10, $\Delta$GEA and $\Delta$Prob rise sharply, showing large gains from higher allowed budgets. Beyond a (maximally allowed) budget of 10, $\Delta$GEA and $\Delta$Prob both stabilize, indicating no further gains. The underlying reason is that the used budget stop increasing after 10 (maximally allowed) when we set the training epochs as 100. In Tables 1 and 2, we report the results with a maximum budget equal to 15.
Attack Training Epochs vs Attack Performance.
Figure 7: Sensitivity analysis w.r.t. maximal training epochs on Syn2 using GXAttack. We observe a steady increase in the used budget as epochs increase. However $\Delta$Prob first increases and then decreases when increasing the training epochs. Similarly, $\Delta$GEA first increases and then decreases with the increase of training epochs. For computation time reason, we only report the results with 100 epochs in Tables 1 and 2. However, it is important to note that the attack performance (in terms of $\Delta$GEA) can be largely improved if we increase the number of training epochs. For instance, $\Delta$GEA can be improved from 0.269 to 0.364 if we increase the training epochs from 100 to 300.
Appendix J Limitations
Our current work suffers from the the following limitations and we plan to mitigate them in the future:
  • GXAttack suffers from low scalability (large requirement of memory), which however can be mitigated by using the Projected Randomized Block Coordinate Descent (PR-BCD) (Geisler et al., 2021);
  • We only employed synthetic datasets. While real-world datasets can provide quantitative evaluation results using metrics like cosine similarity (to measure the consistency of explanations before and after attacks), these metrics may not accurately reflect the true effectiveness of the attacks (i.e., in terms of $\Delta$GEA). This discrepancy suggests that quantitative results can be challenging to interpret, and relying solely on such metrics might not yield meaningful insights. On the other hand, qualitative evaluations are also possible, but as machine learning researchers, our expertise lies not in the realm of such interpretations;
  • GXAttack requires that the target GNN explainer to be differentiable, which is not always the case; However, the attack transferability shows its potential effectiveness on other GNN explainers;
  • We (implicitly) impose an assumption that a small $L_{0}$ perturbation that is prediction-preserving also preserves the causal reasoning for the GNN prediction. Unlike images or texts, it is not easy to verify for graph data. However, we can leverage visualisations to help understand the reasoning process: Figures 8 and  9 give an example which shows the edge perturbations and the corresponding explanation changes after the attacks of GXAttack.
Figure 8: Comparison of original and perturbed graph structures for the target node ‘19’ on graph Syn1. The explanation accuracy decreases from $0.818$ to $0.539$ while the prediction probabilities change from $[0.676,0.324]$ to $[0.844,0.156]$.
Figure 9: The explanation matrices of node ‘19’ before and after attacks, and the difference matrix (original - perturbed).
Appendix K Broader Impacts
Our findings in this paper have significant implications for the development and deployment of GNNs in a variety of domains, particularly in decision-critical applications. Our research highlights a fundamental vulnerability in current post-hoc explanation methods for GNNs, emphasizing that these methods can be unreliable when subject to adversarial perturbations. This revelation calls for a reassessment of how explanations of GNN decisions are generated and trusted, particularly in sectors where transparency and accountability are paramount, such as healthcare, finance, and autonomous systems.
By introducing the GXAttack, an optimization-based adversarial attack specifically designed to target GNN explanations, this paper sets a new standard in the evaluation of explanation methods. This approach not only sheds light on the robustness of existing techniques but also serves as a catalyst for the development of more resilient GNN explanation methods. Ensuring the reliability of explanations in adversarial environments is crucial, as stakeholders rely on these explanations to make informed decisions and justify the use of GNNs in sensitive and impactful contexts.
Moreover, by making our implementation publicly available, we enhance transparency and facilitate widespread involvement in evaluating and enhancing the robustness of GNN explainers. This openness is anticipated to expedite progress in the field of research dedicated to developing more reliable and interpretable graph learning models. Such advancements are likely to increase trust and acceptance of GNN technologies. The primary aim of our research is to contribute to the development of safer and more comprehensible AI systems, ensuring their ethical and responsible application in critical domains.

### v2

@@
 This calls into question the trustworthiness and practical utility of post-hoc explanation methods for GNNs.
-To be able to attack GNN explanation models, we devise a novel attack method dubbed GXAttack, the first optimization-based adversarial attack method for post-hoc GNN explanations under such settings.
+To be able to attack GNN explanation models, we devise a novel attack method dubbed GXAttack, the first optimization-based adversarial white-box attack method for post-hoc GNN explanations under such settings.
 Due to the devastating effectiveness of our attack, we call for an adversarial evaluation of future GNN explainers to demonstrate their robustness.
-1 Introduction
-Graph Neural Networks (GNNs) have achieved promising results in various learning tasks, including representation learning, node classification, graph classification, link prediction, and anomaly detection (Zhou et al., 2020).
-However, intricate data representations and non-linear transformations render their interpretation and thus understanding their predictions non-trivial (Amara et al., 2022).
-This has led to the introduction of GNN explanation methods, which can improve a model’s transparency and thus increase trust in a GNN model, especially when it is deployed in a decision-critical application (e.g., where fairness, privacy, or safety are important) (Ying et al., 2019).
-Moreover, they can help identify the scenarios where a GNN model may fail (Vu and Thai, 2020).
+For reproducibility, our code is available via GitHub.
+Index Terms:
+Graph Neural Networks, Robustness, Explainable Machine Learning
+I Introduction
+Graph Neural Networks (GNNs) [1] have achieved promising results in various learning tasks, including representation learning [2, 3], node classification [4, 5], link prediction [6, 7], and anomaly detection [8, 9].
+However, intricate data representations and non-linear transformations render their interpretation and thus understanding their predictions non-trivial [10].
+This has led to the introduction of GNN explanation methods, which can improve a model’s transparency and thus increase trust in a GNN model, especially when it is deployed in a decision-critical application (e.g., where fairness, privacy, or safety are important) [11].
+Moreover, they can help identify the scenarios where a GNN model may fail [12].
 Explainability of a GNN model can be obtained by designing a self-explainable GNN model, which usually employs a simple model architecture and few parameters, leading to explainability at the cost of suboptimal prediction accuracy; or by extracting post-hoc explanations that do not influence the internal workings of a GNN model and thus maintains its high prediction accuracy.
-In this paper we only consider post-hoc GNN explanation methods, as many such methods have merged over the past decade, including but not limited to Ying et al.
-(2019); Pope et al.
-(2019); Yuan et al.
-(2020); Luo et al.
-(2020); Vu and Thai (2020); Bajaj et al.
-(2021); Yuan et al.
-(2021); Lucic et al.
-(2022); Huang et al.
-(2022).
-It has been shown that traditional deep neural networks (DNNs) are vulnerable to adversarial attacks (Szegedy et al., 2013; Chakraborty et al., 2018), where an imperceptible well-designed perturbation to input data can lead to substantially different prediction results.
-Recently, it was demonstrated that the explanations of predictions made by these DNNs are also fragile (Baniecki and Biecek, 2024).
+In this paper we only consider post-hoc GNN explanation methods, as many such methods have merged over the past decade, including but not limited to [11, 13, 14, 15, 12, 16, 17, 18, 19].
+It has been shown that traditional deep neural networks (DNNs) are vulnerable to adversarial attacks [20, 21], where an imperceptible well-designed perturbation to input data can lead to substantially different prediction results.
+Recently, it was demonstrated that the explanations of predictions made by these DNNs are also fragile [22].
 That is, an unnoticeable, well-designed perturbation to input data can result in completely different explanations while the model’s predictions remain unchanged.
 However, the vulnerability of explanation methods has been primarily studied for DNN models designed for image and text data.
-In contrast, the vulnerability of GNN explanation methods to adversarial attacks has received very limited attention (Fang et al., 2024; Wang et al., 2024), although a plethora of GNN explanation methods have recently been proposed (Yuan et al., 2022; Li et al., 2022; Kakkad et al., 2023).
+In contrast, the vulnerability of GNN explanation methods to adversarial attacks has received very limited attention [23, 24], although a plethora of GNN explanation methods have recently been proposed [25, 26, 27].
 Importantly, existing studies have only investigated the (in)stability of GNN explanations under relatively “mild” settings: they only consider random perturbations and/or prediction-altering perturbations.
-Specifically, Agarwal et al.
-(2022) theoretically analyzed the stability of three GNN explainers under random perturbations.
-Further, Wang et al.
-(2024) considered two scenarios: 1) graphs whose predictions are not changed after randomly flipping $20\%$ edges; and 2) severe perturbations using an off-the-shelf attack algorithm to perturb at most $10\%$ edges to change the GNN predictions (as well as the explanations).
-Besides, Fan et al.
-(2023) proposed an adversarial attack to simultaneously change the GNN predictions and their explanations.
+Specifically, [28] theoretically analyzed the stability of three GNN explainers under random perturbations.
+Further, [24] considered two scenarios: 1) graphs whose predictions are not changed after randomly flipping $20\%$ edges; and 2) severe perturbations using an off-the-shelf attack algorithm to perturb at most $10\%$ edges to change the GNN predictions (as well as the explanations).
+Besides, [29] proposed an adversarial attack to simultaneously change the GNN predictions and their explanations.
 By altering predictions and explanations, however, the graph can become inherently different.
-Fang et al.
-(2024) explored evaluation metrics for GNN explanations from the perspective of adversarial robustness and resistance to out-of-distribution input, where they defined “adversarial robustness" as the difficulty of reversing a GNN prediction by perturbing the complementary of the explanatory subgraph.
-Figure 1: When using post-hoc GNN explainers, the explanatory subgraph on a graph with “prediction-preserving" perturbations (right) can strongly differ from that on the original graph (left).
+[23] explored evaluation metrics for GNN explanations from the perspective of adversarial robustness and resistance to out-of-distribution input, where they defined “adversarial robustness” as the difficulty of reversing a GNN prediction by perturbing the complementary of the explanatory subgraph.
+Figure 1: When using post-hoc GNN explainers, the explanatory subgraph on a graph with “prediction-preserving” perturbations (right) can strongly differ from that on the original graph (left).
 Different from previous studies, we aim to perform perturbations that are both carefully crafted (i.e., optimized rather than random) and imperceptible (i.e., such that the GNN predictions remain unchanged but their explanations differ substantially).
 As shown in Figure 1, we empirically found that small prediction-preserving perturbations can result in largely different explanations generated by post-hoc GNN explainers.
-To our knowledge, we are the first to formally formulate and systematically investigate this problem.
-Specifically, we devise GXAttack, the first optimization-based adversarial attack on post-hoc GNN explanations under this setting.
-We employ a widely used GNN explainer, PGExplainer (Luo et al., 2020), as example target when designing our attack algorithm.
+To our knowledge, we are the first to formally formulate and systematically investigate this problem under a white-box setting.
+Specifically, we devise GXAttack, the first optimization-based adversarial white-box attack on post-hoc GNN explanations.
+We employ a widely used GNN explainer, PGExplainer [15], as example target when designing our attack algorithm.
 Results on various datasets demonstrate the effectiveness of our approach.
-Moreover, our experiments show that other widely used GNN explanation methods, such as GradCAM (Pope et al., 2019), GNNExplainer (Ying et al., 2019), and SubgraphX (Yuan et al., 2021), are also fragile under the attacks optimized for PGExplainer.
+Moreover, our experiments show that other widely used GNN explanation methods, such as GradCAM [13], GNNExplainer [11], and SubgraphX [17], are also fragile under the attacks optimized for PGExplainer.
 Due to the devastating effectiveness of our attack, we call for an adversarial evaluation of future GNN explainers to demonstrate their robustness.
-2 Preliminaries
+We first summarize related work in Section II.
+Following this, we provide necessary background in Section III.
+Then, we formalize the GNN explanation attack problem and introduce our novel attack algorithm GXAttack in Section IV.
+Next, we describe the experimental setups in Section V.
+Experiment results and corresponding analyses are given in Section VI.
+Finally, we conclude the paper in Section VII.
+II Related Work
+II-A GNN Explanations
+[25, 26, 27] perform comprehensive surveys and propose excellent taxonomies of GNN explanation methods.
+Specifically, GNN explanations methods mainly include: 1) Gradients-based methods: Grad [30], GradCAM [13], GuidedBP [31], Integrated Gradients (IG) [32]; 2) Decomposition-based methods: GraphLRP [33], GNN-LRP [34]; 3) Surrogate-based methods: GraphLIME [19], PGM-Explainer [12]; 4) Generation-based methods: XGNN [14]; 5) Perturbation-based methods: GNNExplainer [11], PGExplainer [15], GraphMask [35], SubgraphX [17]; and 6) CF-based methods: CF-GNNExplainer [18], RCExplainer [16].
+II-B Attacks and Defenses of Traditional XAI Methods
+It has been shown that traditional XAI methods (namely those designed to explain deep neural networks for image and text data) are susceptible to malicious perturbations [36, 37, 38].
+Please refer to [22] for a comprehensive survey.
+II-C Robustness/Stability of GNN Explanations
+[25, 28] point out that a reliable GNN explainer should exhibit stability, namely minor perturbations that do not impact the model predictions should not largely change the explanations.
+Particularly, [28] theoretically analyze the stability of three GNN explainers, including vanilla Grad, GraphMask and GraphLIME, for which they derive upper bounds on the instability of explanations.
+Importantluy, [16, 23, 24, 29] are closely related but are different from our work and they are detailed as follows.
+OAR.
+[23] present a novel metric for evaluating GNN explanations, termed Out-Of-Distribution-resistant Adversarial Robustness (OAR).
+OAR aims to solve the limitations of current removal- and generative-based evaluations.
+More concretely, they evaluate post-hoc explanation subgraphs by computing their robustness under attack, which consists of three steps: 1) they formulate the adversarial robustness tailored for GNN explanations problem, which is the minimum adversarial perturbation on the structure of complementary subgraph; 2) they introduce a tractable and easy-to-implement objective of adversarial robustness for GNN explanations; after relaxation, the objective becomes a constrained graph generation problem (namely random perturbations) rather than an adversarial attack problem; 3) they present an Out-Of-Distribution (OOD) reweighting block that aims to confine the evaluation on original data distribution.
+This work is an initial attempt to explore evaluation metrics for GNN explanations from the perspectives of adversarial robustness and resistance to OOD.
+In contrast, we focus on optimization-based adversarial attack on post-hoc GNN explanations rather than providing an evaluation metric.
+V-InFoR.
+[24] point out that “existing GNN explainers are not robust to the structurally corrupted graphs, namely graphs with noisy or adversarial edges.” This study investigates the negative effect of minor structural corruptions (which do not change the predictions) and severe structural corruptions (which change the predictions) on GNN explainers.
+Particularly, they provide quantitative evidence that existing GNN explainers are fragile to structurally corrupted graphs.
+They evaluate 6 GNN explainers under two settings: 1) minor corruptions where they select the perturbed graphs whose predictions are not changed after randomly flipping 20% edges; 2) severe corruptions where they employ adversarial attack algorithm GRABNEL [39] to perturb at most 10% edges to change the predictions.
+In contrast, our method focus on optimization-based (rather than random) and prediction-preserving (rather than prediction-altering) adversarial attacks.
+RCExplainer.
+[16] aim to find a small set of edges of the input graph such that the prediction result will substantially change if we remove those edges.
+Specifically, they propose RCExplainer to generate robust counter-factual explanations for GNNs in two steps: 1) they employ a set of decision regions to model the decision logic of a GNN, where each decision region determines the predictions on multiple graphs that are predicted to be the same class; 2) they leverage a DNN model to explore the decision logic, and thereby extract robust counter-factual explanations as a small set of edges of the input graph.
+This method is only applicable to GNNs that belong to Piecewise Linear Neural Networks.
+In this paper, we focus on attacking differentiable post-hoc GNN explanation methods that can be applied to any GNNs.
+GEAttack.
+[29] empirically demonstrate that GNN explanation methods can be employed as tools to detect the adversarial perturbations on graphs.
+On this basis, they propose an attack framework dubbed GEAttack that can jointly attack both GNN and its explanations.
+Specifically, they formulate GEAttack as a bi-level optimization problem: 1) they mimic the GNN explanation method optimization process to obtain graph explanations in the inner loop; 2) they compute the gradient of the attack objective w.r.t.
+the explanations in the outer loop.
+Different from GEAttack, we consider prediction-preserving attacks, which are inherently more challenging.
+Concurrent Work.
+While the concurrent work [40] also addresses a similar topic, our study provides an alternative perspective by performing white-box attacks (namely we assume knowing the full knowledge about the GNN classifier and the GNN explainer), while they perform gray-box attacks by assuming knowing only the explanation loss and the generated explanatory edges of the GNN explainer.
+II-D Adversarial Robustness of GNNs
+Starting with the seminal works [41, 42], that both proposed adversarial attacks on GNNs, a rich literature formed in the realm of GNN-specific adversarial attacks, defenses, and certifications [43, 44].
+Most attacks primarily focus on altering predictions via perturbations of the discrete graph structure (edge insertions/deletions), either between existing nodes [41] or via the insertion of new (adversarial) nodes [45].
+Optimization-based attacks either operate globally (number of edge insertions/deletions) [46, 47], alter a local neighborhood [41], or can handle a combination of both constraints [48].
+While adversarial attacks on GNN explanation methods may benefit from optimization methods over the discrete graph structure, a careful discussion and derivation of appropriate attack objectives and metrics was missing.
+III Preliminaries
 We utilize lowercase letters, bold lowercase letters, uppercase letters and calligraphic fonts to represent scalars ($x$), vectors ($\mathbf{x}$), matrices ($\mathbf{X}$), and sets ($\mathcal{X}$), respectively.
@@
 Thus, a graph can also be represented as $\mathcal{G}=(\mathbf{A},\mathbf{X})$.
-2.1 Node classification with GNNs
+III-A Node Classification with GNNs
 We consider the task of node classification.
@@
 In other words, we have $f_{\theta}(\mathcal{G}(v))=c$, where $\mathcal{G}(v)$ denotes the computation graph of node $v$.
-Typically, $f_{\theta}(\cdot)$ consists of 1) a GNN model used to learn node representations, and 2) a multi-layer perceptron (MLP) classifier that assigns labels.
+Typically, $f_{\theta}(\cdot)$ consists of two components: 1) a GNN model used to learn node representations, and 2) a multi-layer perceptron (MLP) classifier that assigns labels.
 More specifically, given a set of training nodes and their labels $\{(v_{1},y_{1}),...,(v_{n},y_{n})\}$ from graph $\mathcal{G}$, the objective function of a GNN classifier can be defined as
 […]
+$=\sum_{i=1}^{n}\sum_{c}^{C}\mathbb{I}(y_{i}=c)\ln(f_{\theta}(\mathcal{G}(v_{i}),y_{i})^{(c)}),$ (2)
 where $\theta$ is the set of learnable parameters for $f_{\theta}(\cdot)$, $l(\cdot)$ is the cross-entropy loss function, and $\mathbb{I}(\cdot)$ denotes the Kronecker function.
@@
 $f_{\theta}(\mathbf{A},\mathbf{X}):=\operatorname{softmax}(\tilde{\mathbf{A}}\sigma(\tilde{\mathbf{A}}\mathbf{X}\mathbf{W}_{1})\mathbf{W}_{2}).$ (4)
-2.2 GNN Explainer
+III-B GNN Explainer
 We consider post-hoc GNN explanation methods that provide instance-level explanations: given a graph $\mathcal{G}$, a target node $v$, and prediction $f_{\theta}(\mathcal{G}(v))$, the explainer $g_{\lambda}(v,f(\mathcal{G}(v)))$ aims to ‘explain’ this particular prediction.
-Following common explanation methods, we assume that explanations are given in the form of an explanatory subgraph, namely $\mathcal{G}_{s}(v)=g(v,f(\mathcal{G}(v)))$, consisting of a small set of important edges and their associated nodes, where the explanatory subgraph is usually connected and edge importance is measured by a so-called edge importance score.
-In this paper, we take PGExplainer (Luo et al., 2020) as an example to attack for three reasons.
-First, Faber et al.
-(2021) pointed out that GNN explanations providing edge importance are preferable to explanations providing node importance, due to the fact that edges have more fine-grained information than nodes.
-Second, Agarwal et al.
-(2023) empirically showed that PGExplainer generates the least unstable explanations, exhibiting a $35.35\%$ reduction in explanation instability relative to the average instability of other GNN explainers.
+Following common explanation methods, we assume that explanations are given in the form of an explanatory subgraph, namely $\mathcal{G}_{s}(v)=g(v,f(\mathcal{G}(v)))$, consisting of a small set of important edges and their associated nodes, where the edge importance is measured by a so-called edge importance score.
+In this paper, we take PGExplainer [15] as an example to attack for three reasons.
+First, [49] pointed out that GNN explanations providing edge importance are preferable to explanations providing node importance, due to the fact that edges have more fine-grained information than nodes.
+Second, [50] empirically showed that PGExplainer generates the least unstable explanations, exhibiting a $35.35\%$ reduction in explanation instability relative to the average instability of other GNN explainers.
 Third, the objective function of PGExplainer is differentiable.
-For completeness, we briefly revisit the goal of PGExplainer (see Appendix A for more details).
+For completeness, we briefly revisit the goal of PGExplainer.
 For brevity, we denote $\mathcal{G}(v)$ as $\mathcal{G}$ (and $\mathcal{G}_{s}(v)$ as $\mathcal{G}_{s}$ analogously) throughout the remainder of the paper.
@@
 where $H(\mathbf{y})$ is the entropy of $f_{\theta}(\mathcal{G})$, and $H(\mathbf{y}|\mathcal{G}_{s})$ is the conditional entropy of the GNN’s prediction on the explanatory graph, namely $f_{\theta}(\mathcal{G}_{s})$.
-3 GXAttack: GNN Explanation Adversarial Attacks
-Given an input $\mathcal{G}$, a GNN classifier $f_{\theta}(\cdot)$, and a GNN explainer $g_{\lambda}(\cdot)$, the original prediction for the target node $v$ is $f_{\theta}(\mathcal{G}(v))$ and its corresponding original explanatory subgraph is given by $\mathcal{G}_{s}(v)=g_{\lambda}(f_{\theta}(v,\mathcal{G}(v)))$.
+IV GXAttack:
+GNN Explanation Adversarial Attacks
+Given an input graph $\mathcal{G}$, a GNN classifier $f_{\theta}(\cdot)$, and a GNN explainer $g_{\lambda}(\cdot)$, the original prediction for the target node $v$ is $f_{\theta}(\mathcal{G}(v))$ and its corresponding original explanatory subgraph is given by $\mathcal{G}_{s}(v)=g_{\lambda}(f_{\theta}(v,\mathcal{G}(v)))$.
 From an attacker’s perspective, a desirable manipulated graph $\mathcal{G}_{adv}:=\hat{\mathcal{G}}=\phi({\mathcal{G}})$ for attacking node $v$ should have the following properties, where $\phi(\cdot)$ is an attack function that perturbs its input, i.e., it removes and/or adds a small set of edges:
@@
     The explanatory subgraph for node $v$ on original graph $\mathcal{G}_{s}(v)=g_{\lambda}(f_{\theta}(v,\mathcal{G}(v)))$ and that on perturbed graph $\hat{\mathcal{G}}_{s}(v)=g_{\lambda}(f_{\theta}(\hat{\mathcal{G}}(v)))$ are substantially different: $\mathcal{G}_{s}(v)\neq\hat{\mathcal{G}}_{s}(v)$.
-In particular, we aim to make $\hat{\mathbf{A}}_{s}$ strongly different from $\mathbf{A}_{s}$.
-3.1 Attack Objectives
+In particular, we aim to make $\hat{\mathbf{A}}_{s}$ (i.e., the adjacency matrix of explanatory subgraph on perturbed graph) strongly different from $\mathbf{A}_{s}$ (i.e., the adjacency matrix of explanatory subgraph on original graph).
+IV-A Attack Objectives
 Generic Objective. This paper focuses on adversarial structure attacks on post-hoc GNN explanation for node classification:
@@
 Loss Function. We should instantiate the loss function $\mathcal{L}$ such that it pertains to the desirable properties of manipulations discussed before. More specifically, it is defined as
-$\mathcal{L}:=\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))).$ (6)
+$\mathcal{L}:=~{}$ $\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))$
+$-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\hat{\mathbf{A}},\mathbf{X}))).$ (7)
 The first term ensures that the GNN predictions before and after perturbations are approximately the same.
@@
 We directly model the edge perturbation matrix $\mathbf{P}\in\{0,1\}^{n\times n}$, where we flip $\mathbf{A}_{ij}$ (namely $0\to 1$ or $1\to 0$) if $\mathbf{P}_{ij}=1$, and keep $\mathbf{A}_{ij}$ intact otherwise.
-On this basis, we have
-$\underset{\mathbf{P}\in\{0,1\}^{n\times n}\text{~{}s.t.~{}}\sum\mathbf{P}_{ij}<B}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))),$ (7)
+On this basis, Eq 6 can be rewritten as
+$\underset{\mathbf{P}\in\{0,1\}^{n\times n}\text{~{}s.t.~{}}\sum\mathbf{P}_{ij}<B}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))$
+$-\beta*\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\mathbf{P},\mathbf{X}))),$ (8)
 where $\oplus$ means element-wise exclusive or operation, and $B$ is the edge flipping budget.
-3.2 Relaxations and Loss Definitions
+IV-B Relaxations and Loss Definitions
 Unfortunately, directly minimizing $\mathcal{L}$ w.r.t.
@@
 Next, we instantiate $P(m_{ij})$ using a Bernoulli distribution, namely $P(m_{ij})\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{ij})$ with $P(m_{ij}=1)=\tilde{\mathbf{P}}_{ij}$.
-As a result, Eq 7 is equivalent to
+As a result, Eq IV-A is equivalent to
 $\underset{\mathbf{P}}{{\max}}~{}\mathcal{L}(\mathbf{P})=\underset{\mathcal{M}}{\max}~{}\mathbb{E}_{\mathcal{M}}[\mathcal{L}(\mathcal{M})]\approx\underset{\tilde{\mathbf{P}}}{\max}~{}\mathbb{E}_{\mathcal{M}\sim\tilde{\mathbf{P}}}[\mathcal{L}(\mathcal{M})].$ (9)
@@
 Hence, we can utilise gradient-based methods to optimize the objective function.
-At the last step of the attack, we then discretize the result (see Section 3.3).
+At the last step of the attack, we then discretize the result (see Section IV-C).
 Loss Term 1. We approximate the first term in $\mathcal{L}$ as follows:
-$\underset{\tilde{\mathbf{P}}}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X})):=\underset{\tilde{\mathbf{P}}}{\max}~{}H(f_{\theta}(\mathbf{A},\mathbf{X}))-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))\$ (9)
-$\propto\underset{\tilde{\mathbf{P}}}{\max}~{}-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))\approx\underset{\tilde{\mathbf{P}}}{\max}~{}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}})).$ (10)
-Eq 9 holds by definition of mutual information; Eq 10 left half holds as $\theta,\mathbf{A},\mathbf{X}$ are fixed; Eq 10 right half holds when we replace conditional entropy $H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ with cross entropy $H(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$.
-We define $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}+\tilde{\mathbf{P}}_{ij}$ if $\mathbf{A}_{ij}=0$, and $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}-\tilde{\mathbf{P}}_{ij}$ otherwise.
+$\underset{\tilde{\mathbf{P}}}{\max}~{}\text{MI}(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ (10)
+$:=\underset{\tilde{\mathbf{P}}}{\max}~{}H(f_{\theta}(\mathbf{A},\mathbf{X}))-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ (11)
+$\propto\underset{\tilde{\mathbf{P}}}{\max}~{}-H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ (12)
+$\approx\underset{\tilde{\mathbf{P}}}{\max}~{}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}})).$ (13)
+Eq 11 holds by definition of mutual information; Eq 12 holds as $\theta,\mathbf{A},\mathbf{X}$ are fixed; Eq 13 holds when we replace conditional entropy $H(f_{\theta}(\mathbf{A},\mathbf{X})|f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ with cross entropy $H(f_{\theta}(\mathbf{A},\mathbf{X}),f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X}))$ for practical optimization reason.
+Moreover, we define $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}+\tilde{\mathbf{P}}_{ij}$ if $\mathbf{A}_{ij}=0$, and $\mathbf{A}_{ij}\oplus\tilde{\mathbf{P}}_{ij}$ = $\mathbf{A}_{ij}-\tilde{\mathbf{P}}_{ij}$ otherwise.
 Loss Term 2. Next, we approximate the second term in $\mathcal{L}$ as follows:
-$\underset{\tilde{\mathbf{P}}}{\min}~{}\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X})))$ $\propto\underset{\tilde{\mathbf{P}}}{\max}~{}\text{Dist}(\mathbf{M}_{s},\hat{\mathbf{M}}_{s})$ (11)
+$\underset{\tilde{\mathbf{P}}}{\min}~{}\text{MI}(g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X})),g_{\lambda}(f_{\theta}(\mathbf{A}\oplus\tilde{\mathbf{P}},\mathbf{X})))$ (14)
+$\propto\underset{\tilde{\mathbf{P}}}{\max}~{}\text{Dist}(\mathbf{M}_{s},\hat{\mathbf{M}}_{s})$ (15)
 $=:\underset{\tilde{\mathbf{P}}}{\max}~{}\left(1-\frac{\text{vec}(\mathbf{M}_{pr})\cdot\text{vec}(\hat{\mathbf{M}}_{pr})}{\|\mathbf{M}_{pr}\|_{F}\|\hat{\mathbf{M}}_{pr}\|_{F}}\right),$ (16)
+where $\mathbf{M}_{s}$ and $\hat{\mathbf{M}}_{s}$ are the explanatory subgraphs on original graph and perturbed graph, respectively; $\mathbf{M}_{pr}$ and $\hat{\mathbf{M}}_{pr}$ are the predicted explanations (edge importance matrices) on original graph and perturbed graph, respectively.
+In other words, the explanatory subgraph is represented as an edge importance matrix in this study.
 Eq 15 holds by definition; Eq 16 holds when we define the $Dist$ function as the cosine distance metric, where $\text{vec}(\cdot)$ flattens a matrix to a vector, and $\|\cdot\|_{F}$ represents the Frobenius Norm.
-Final Relaxed Objective. As a result, optimizing Eq 7 amounts to optimizing
-$\underset{\tilde{\mathbf{P}}\in[0,1]^{n\times n}\text{~{}s.t.~{}}\sum\tilde{\mathbf{P}}_{ij}<B}{\max}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}}))+\beta*\left(1-\frac{\text{vec}(\mathbf{M}_{pr})\cdot\text{vec}(\hat{\mathbf{M}}_{pr})}{\|\mathbf{M}_{pr}\|_{F}\|\hat{\mathbf{M}}_{pr}\|_{F}}\right),$ (13)
-where $\mathcal{L}(\tilde{\mathbf{P}})$ is differentiable w.r.t. $\tilde{\mathbf{P}}$.
-3.3 Optimization
-We employ Projected Gradient Ascent Algorithm (based on Xu et al.
-(2019)) to optimize objective function (13), and the details are given in Algorithm 1.
+Final Relaxed Objective. As a result, optimizing Eq IV-A amounts to optimizing
+$\underset{\tilde{\mathbf{P}}\in[0,1]^{n\times n}\text{~{}s.t.~{}}\sum\tilde{\mathbf{P}}_{ij}<B}{\max}$ $\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\hat{\mathcal{G}}))$
+$+\beta*\left(1-\frac{\text{vec}(\mathbf{M}_{pr})\cdot\text{vec}(\hat{\mathbf{M}}_{pr})}{\|\mathbf{M}_{pr}\|_{F}\|\hat{\mathbf{M}}_{pr}\|_{F}}\right).$ (17)
+As a result, $\mathcal{L}(\tilde{\mathbf{P}})$ (i.e., Eq IV-B) is differentiable w.r.t. $\tilde{\mathbf{P}}$.
+IV-C Optimization
+We employ Projected Gradient Ascent Algorithm (based on [46]) to optimize objective function (IV-B), and the details are given in Algorithm 1.
 To go back from the continuous $\tilde{\mathbf{P}}_{ij}$ to discrete ${\mathbf{P}}_{ij}$, we perform Bernoulli sampling with ${\mathbf{P}}_{ij}\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{ij})$ after optimization (Line 10).
-We note that the Projected Randomized Block Coordinate Descent (PR-BCD) (Geisler et al., 2021) can perform this optimization more efficiently, but do not compare both methods due to the low scalability of the evaluated explainers.
-The complexity analysis of GXAttack is given in Appendix C.
+We note that the Projected Randomized Block Coordinate Descent (PR-BCD) [47] can perform this optimization more efficiently, but do not compare both methods due to the low scalability of the evaluated explainers.
 Algorithm 1 Pseudo-code for GXAttack
@@
 9:    $\tilde{\mathbf{P}}_{t}\leftarrow\tilde{\mathbf{P}}_{t-1}+\alpha_{t}\cdot\nabla_{\tilde{\mathbf{P}}_{t-1}}\mathcal{L}(\hat{\mathbf{y}},\mathbf{y},\hat{\mathcal{G}}_{s},\mathcal{G}_{s})$triangleright Update gradients
-10:    $\tilde{\mathbf{P}}_{t}\leftarrow\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})\leq B]}(\tilde{\mathbf{P}}_{t})$triangleright Project gradients, see Appendix B for more detail
+10:    $\tilde{\mathbf{P}}_{t}\leftarrow\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})\leq B]}(\tilde{\mathbf{P}}_{t})$triangleright Project gradients, see Appendix Appendix A: Projected Gradient Descent for more detail
 11:end for
@@
 13:Return $\mathbf{A}\oplus\mathbf{P}$triangleright Which is defined as $\mathbf{A}_{ij}+\mathbf{P}_{ij}(1-2\mathbf{A}_{ij})$
-4 Experimental Setup
+IV-D Complexity Analysis
+We first discuss the dense case here and then make a discussion of GNNs/explainers implemented with sparse matrices. The time complexity is as follows:
+  1. 1.
+    Initial Operations: $O(f_{\theta}+g_{\lambda})$, typically $O(n^{2})$ for a dense GNN and explainer (sometimes even $O(n^{3})$).
+  2. 2.
+    Each Epoch:
+    • Prediction and explanation on perturbed graph: $O(2T\cdot(f_{\theta}+g_{\lambda}))$, where $T$ is the number of epochs.
+    • Gradient computation and update: Dependent on the efficiency of backpropagation through the GNN, but generally $O(T\cdot n^{2})$ for gradient computation and a similar order for updates.
+  3. 3.
+    Sampling Step: $O(n^{2})$.
+Meanwhile, the space complexity is as follows:
+  1. 1.
+    Space for Storing Matrices: Storing $\mathbf{A}$, $\mathbf{X}$, $\tilde{\mathbf{P}}$ each requires $O(n^{2})$ space.
+  2. 2.
+    Additional Space for Operations: Space required for intermediate gradients and outputs during predictions, typically $O(n^{2})$.
+Since the explainer uses dense matrices, we also choose an attack operating on all $n^{2}$ edges.
+Thus, the total complexity $O((T+1)\cdot(f_{\theta}+g_{\lambda})+n^{2})$ since we have $T$ attack steps plus the initial predictions as well as explanations and, last, we sample the final perturbations.
+Assuming that $O(f_{\theta})=O(g_{\lambda})=O(n^{2}d+nd^{2})$ and $d\ll n$, this yields a total time complexity of $O(n^{2})$ with number of attack iterations $T\ll n$, and number of nodes $n$.
+This is the same asymptotic complexity as the dense GNN $f_{\theta}$ and explainer $g_{\lambda}$.
+Assuming GNN and explainer are in $O(f_{\theta})=O(g_{\lambda})=O(Ed+nd^{2})$ with number of edges $E$, a dense attack would dominate the overall complexity ($O(n^{2})$).
+An immediate remedy would yield PRBCD [47] since its complexity is $O(B+E)$ with edge flipping budget $B$ and all additional components (e.g., loss function) could be implemented in $O(E)$ with sparse matrices.
+This yields an overall time and space complexity of $O(E)$, assuming $B\ll E$ and hidden dimensions $d\ll n$, which is the same asymptotic complexity as the sparse GNN $f_{\theta}$ and explainer $g_{\lambda}$.
+V Experimental Setup
 We aim to answer the following research questions: (RQ1) How effective are the attacks generated by GXAttack on PGExplainer?
@@
 and (RQ3) Can the attacks generated by GXAttack (optimized for PGExplainer) transfer to other GNN explanation methods?
-4.1 Datasets
+V-A Datasets
 To answer these questions, we need to evaluate the GNN explanation quality before and after the attacks.
-However, as pointed out by Agarwal et al.
-(2023), evaluating the quality of post-hoc GNN explanations is challenging.
+However, as pointed out by [50], evaluating the quality of post-hoc GNN explanations is challenging.
 This is because existing benchmark graph datasets lack ground-truth explanations or their explanations are not reliable.
@@
 As a result, ShapeGGen allows us to mimic graph data in various real-world applications.
-More importantly, ShapeGGen can ensure that the generated data and its ground-truth explanations do not suffer from GNN explanation evaluation pitfalls (Faber et al., 2021), namely trivial explanations, redundant explanations, and weak GNN predictors.
+More importantly, ShapeGGen can ensure that the generated data and its ground-truth explanations do not suffer from GNN explanation evaluation pitfalls [49], namely trivial explanations, redundant explanations, and weak GNN predictors.
 This makes ShapeGGen very suitable for studying the limitations of GNN explainers.
-We consider the synthetic datasets in Table 3 (See Appendix D).
+We consider the synthetic datasets in Table I.
+TABLE I: Summary of synthetic datasets with ground-truth explanations.
+#Nodes and #Edges represent the number of expected nodes and edges, respectively.
+We do not consider larger datasets for two reasons: 1) explanation accuracy will be very low ($<0.3$) for most explainers, and it is not meaningful to attack low accuracy explainers; 2) computation will be slow for the attacker (and the explainer).
+Dataset Syn1 Syn2 Syn3 Syn4 Syn5 Syn6 Syn7
+Shape House House House Circle House House House
+#Subgraphs 10 50 50 50 50 50 100
+Subgraph Size 6 6 10 10 10 10 10
+P(Connection) 0.3 0.06 0.06 0.06 0.12 0.20 0.06
+#Classes 2 2 2 2 2 2 2
+#Nodes 59 299 521 511 489 492 1018
+#Edges 164 970 1432 1314 1664 1898 3374
+Average Degree 2.8 3.2 2.7 2.6 3.4 3.9 3.3
+Max Budget 10 15 15 15 15 15 15
 Why not real-world datasets?
 We do not consider real-world datasets for two key reasons: 1) synthetic datasets with ground truth are de-facto standards for evaluating post-hoc GNN explainers (for node classification), including GNNExplainer, PGExplainer, PGM-Explainer, SubgraphX, CF-GNNExplainer, RCExplainer, etc.
-(see Table 11 for more details); 2) without ground-truth information we can only compute some metrics, such as cosine similarity of explanation before and after attack (which will be defined later) to quantify attack performance.
-However, as shown in Table 1, the cosine similarity and the real attack performance (in terms of $\Delta\text{GEA}$, which will be defined later) are not necessarily related.
+2) without ground-truth information we can only compute some metrics, such as cosine similarity of explanation before and after attack (which will be defined later) to quantify attack performance.
+However, as shown in Table II, the cosine similarity and the real attack performance (in terms of $\Delta\text{GEA}$, which will be defined later) are not necessarily related.
 Therefore, considering such metrics may not be meaningful.
-4.2 Metrics and Baselines
+V-B Metrics and Baselines
 Graph Explanation Accuracy (GEA).
-We employ the graph explanation accuracy proposed in Agarwal et al.
-(2023) to quantify the quality of generated explanations when the ground-truth explanations are available.
+We employ the graph explanation accuracy proposed in [50] to quantify the quality of generated explanations when the ground-truth explanations are available.
 For simplicity, we assume that each edge is binary coded as ‘0’ (non-contributory) or ‘1’ (contributory) to model predictions in both ground-truth (provided by ShapeGGen) and predicted explanation masks (provided by the GNN explainer).
 Specifically, graph explanation accuracy measures the correctness of generated explanations by computing the Jaccard Index between ground-truth explanation $\mathbf{M}_{gt}$ and predicted explanation $\mathbf{M}_{pr}$ as
-$\text{JAC}(\mathbf{M}_{gt},\mathbf{M}_{pr})=\frac{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})}{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FN}(\mathbf{M}_{gt},\mathbf{M}_{pr})},$ (14)
+$\text{JAC}(\mathbf{M}_{gt},\mathbf{M}_{pr})$
+$=\frac{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})}{\text{TP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FP}(\mathbf{M}_{gt},\mathbf{M}_{pr})+\text{FN}(\mathbf{M}_{gt},\mathbf{M}_{pr})},$ (18)
 where TP, FP, FN denote True Positives, False Positives, and False Negatives, respectively.
@@
 where $\text{vec}(\cdot)$ flattens its input as a vector, and $\|\cdot\|_{F}$ represents the Frobenius Norm.
-Higher values of cos_sim (within $[-1,1]$) indicate higher graph explanation stability, and thus less successful attacks.
+Higher values of $\operatorname{Sim}_{\cos}$ (within $[-1,1]$) indicate higher graph explanation stability, and thus less successful attacks.
 Prediction Change.
-To measure the impact of edge perturbations on GNN predictions, we consider 1) $\Delta$Label: the ratio of nodes of which the predicted label has changed after attacks; and 2) $\Delta$Prob: the average absolute change of predicted probability (of the original predicted class) of all nodes.
-Baselines.
+To measure the impact of edge perturbations on GNN predictions, we consider two metrics as follows: 1) $\Delta$Label: the ratio of nodes of which the predicted label has changed after attacks; and 2) $\Delta$Prob: the average absolute change of predicted probability (of the original predicted class) of all nodes, which is formally defines as
+$\Delta\text{Prob}=\frac{\sum_{i=1}^{N}\lvert P(y_{i}=y_{i}^{*})-\hat{P}(y_{i}=y_{i}^{*})\rvert}{N},$
+where $P(y_{i}=y_{i}^{*})$ is the prediction probability of node $v_{i}$ being classified as $y_{i}^{*}$ (the original predicted class with the highest probability) on the original graph.
+Moreover, $\hat{P}(y_{i}=y_{i}^{*})$ is the prediction probability of node $v_{i}$ being classified as $y_{i}^{*}$ (the original predicted class with the highest probability on the original graph) on the perturbed graph.
+Attacker Baselines.
 Given the novelty of our problem setting, we only consider two trivial baselines for edge perturbations: 1) random flipping: we randomly flip $B_{1}$ edges; and 2) random rewiring: we randomly rewire $B_{2}$ edges within k-hop neighbours of the target node.
-4.3 Attack Transferability Settings
-We consider attack transferability on the following four types of GNN explainers: 1) Gradient-based: Grad (Simonyan et al., 2013), GradCAM (Pope et al., 2019), GuidedBP (Baldassarre and Azizpour, 2019), Integrated Gradients (IG) (Sundararajan et al., 2017); 2) Perturbation-based: GNNExplainer (Ying et al., 2019), SubgraphX (Yuan et al., 2021); 3) Surrogate-based: PGMExplainer (Vu and Thai, 2020); and 4) Random Explainer that randomly generates random important edges.
-For space reasons, all training protocols are deferred to Appendix G. Code is given in anonymous2024code.
-5 Experimental Results and Analysis
-Before answering the research questions, we state three presumptions and check them empirically in Section 5.1. Next, we answer RQ1 and RQ2 in Section 5.2, and RQ3 in Section 5.3.
-5.1 Experiments to Check Presumptions
+V-C Attack Transferability Settings
+We consider attack transferability on the following five types of GNN explainers: 1) Gradient-based: GradCAM [13], Integrated Gradients (IG) [32]; 2) Perturbation-based: GNNExplainer [11], SubgraphX [17]; 3) Surrogate-based: PGMExplainer [12]; 4) Decomposition-based methods: GNN-LRP [34] and 5) Random Explainer that randomly generates random important edges.
+V-D Training Protocols and Compute Resources
+GNN Predictor Training Protocols.
+To avoid pitfalls caused by weak GNN predictor [49], we should train the GNN predictor close to its maximum possible performance.
+When training the GNN predictor, we employ an Adam optimizer where the learning rate is 1e-2, weight decay is 1e-5, the number of epochs is 300, and the hidden dimension is 32.
+GNN Explainer Training Protocols.
+We followed the authors’ recommendations for setting the hyperparameters of GNN explanation methods.
+We select top-$k~{}(k=25\%)$ important edges to generate explanations for all GNN explainers.
+Particularly, when training the PGExplainer, the number of training epochs is 10.
+GNN Explanation Attacker Training Protocols.
+We set the trade-off hyperparameter in the final loss as $0.1$ such that the magnitudes of the two loss terms are approximately the same.
+Besides, the learning rate in gradient ascent is set to $0.1$.
+Importantly, we set the maximal training epochs as 100 due to computation time reason, while further increasing this number can largely improve the attack performance (in terms of $\Delta$GEA).
+Moreover, we set the perturbation budget to 15 at maximum for synthetic data (note that we attempt to employ a hard minimum budget of 5 with 2000 maximum trials of Bernoulli sampling).
+We perform sensitivity analysis regarding these hyperparameters, and the results and analysis will be discussed in Section VI-D.
+Explanation Accuracy Computation.
+The $get\_acc(\cdot)$ function of GraphXAI considers all non-zero edges as predicted positives (namely with binary code ‘1’) by default when computing JAC.
+Following [50], we set the top 25% (of edges with the highest edge importance scores) as 1 and the rest as zeros in our evaluations when computing both GEA and cosine similarity.
+Software and Hardware.
+The evaluated GNN explainers are implemented mainly based on GraphXAI  [^1] https://github.com/mims-harvard/GraphXAI [50].
+Algorithms are implemented in Python 3.8 (using PyTorch [51] and PyTorch Geometric [52] libraries when applicable) and ran on internal clusters with AMD 128 EPYC7702 CPUs (with 512GB RAM).
+All experiments can be finished within 7 days if using 10 such machine instances.
+All code and datasets are available on GitHub [^2] https://github.com/ZhongLIFR/GXAttack.
+VI Experimental Results and Analysis
+Before answering the research questions, we state three presumptions and check them empirically in Section VI-A.
+Next, we answer RQ1 and RQ2 in Section VI-B, and RQ3 in Section VI-C.
+Following this, we perform sensitivity analysis in Section VI-D and property analysis in Section VI-E.
+VI-A Experiments to Check Presumptions
 Presumption 1.
@@
 For each dataset, we first get the set of correctly predicted nodes and the set of wrongly predicted nodes; then we compare the distribution of their explanation accuracy.
-As shown in Figure 2 in Appendix E, there is no notable difference.
-In other words, we do not observe that wrongly labelled nodes tend to have lower explanation accuracy despite Faber et al.
-(2021) claim that it is unfair to use wrongly labelled nodes to evaluate GNN explanations (as they may not use the ground-truth explanations to make predictions and thus tend to have lower explanation accuracy).
+As shown in Figure 6 in Appendix, there is no notable difference.
+In other words, we do not observe that wrongly labelled nodes tend to have lower explanation accuracy despite [49] claim that it is unfair to use wrongly labelled nodes to evaluate GNN explanations (as they may not use the ground-truth explanations to make predictions and thus tend to have lower explanation accuracy).
 Presumption 2.
@@
 For each dataset, we first divide the nodes into five groups based on their prediction confidence; then we compare the distribution of their explanation accuracy.
-Tables 410 in Appendix E show that group ‘G9’ (i.e., the set of nodes with prediction confidence in $[0.9,1.0]$) indeed tends to have lower explanation accuracy on the original graph than other groups.
+Table IV in Appendix show that group ‘G9’ (i.e., the set of nodes with prediction confidence in $[0.9,1.0]$) indeed tends to have lower explanation accuracy on the original graph than other groups.
 Presumption 3.
@@
 To check this presumption, we first divide the nodes into five groups based on their prediction confidence; then we compare the distribution of their attack performance.
-Tables 410 in Appendix E show that group ‘G9’ also tends to have less successful attacks (i.e., smaller $\Delta\text{GEA}$) when compared to other groups.
+Table IV in Appendix show that group ‘G9’ also tends to have less successful attacks (i.e., smaller $\Delta\text{GEA}$) when compared to other groups.
 Possible reasons are that: 1) the nodes in this group have lower explanation accuracy on original graph; and 2) the prediction confidence is close to 1, leaving no much room for perturbations.
 These two facts together make the attack more challenging.
-5.2 Experiments to Answer RQ1 and RQ2: Attack Effectiveness
-Table 1: Results on all synthetic datasets (average $\pm$ standard deviations across 5 trials).
+VI-B Experiments to Answer RQ1 and RQ2: Attack Effectiveness
+TABLE II: Results on all synthetic datasets (average $\pm$ standard deviations across 5 trials).
 O.
@@
 $\Delta$Label is the average change of predicted labels, while $\Delta$Prob denotes the average absolute change of predicted probability (of the original predicted class).
-“GXAttack", “Rnd.
-Flipping", and “Rnd.
-Rewiring" correspond to our attack method, random flipping, and random rewiring, respectively.
+“GXAttack”, “Rnd.
+Flipping”, and “Rnd.
+Rewiring” correspond to our attack method, random flipping, and random rewiring, respectively.
 $\uparrow^{\prime}$ indicates larger values are preferred while $\downarrow^{\prime}$ means the opposite.
@@
 $\Delta$Prob()($\downarrow$)( ↓ ) $0.234_{\pm\text{0.016}}$ $0.210_{\pm\text{0.013}}$ $0.198_{\pm\text{0.007}}$ $0.200_{\pm\text{0.003}}$ $0.221_{\pm\text{0.005}}$ $0.174_{\pm\text{0.013}}$ $0.259_{\pm\text{0.005}}$
-From Table 1, we have the following main observations:
+From Table II, we have the following main observations:
 Observation 1.
@@
 The type of motifs (namely ground-truth explanations) appears to have minimal impact on both explanation accuracy and attack performance.
-This can be obtained by comparing the results on Syn3 (Motif =“House") with those on Syn4 (Motif =“Circle").
+This can be obtained by comparing the results on Syn3 (Motif =“House”) with those on Syn4 (Motif =“Circle”).
 Observation 3.
@@
 Answer to RQ1.
-From Table 1, we can see that GXAttack can achieve a $\Delta\text{GEA}$ ranging from $0.088$ to $0.267$ while using a very small perturbation budget (less than 7.1 on all datasets) and keeping the predicted labels nearly unchanged (i.e., $\Delta\text{Label}$ is nearly zero) after perturbation.
+From Table II, we can see that GXAttack can achieve a $\Delta\text{GEA}$ ranging from $0.088$ to $0.267$ while using a very small perturbation budget (less than 7.1 on all datasets) and keeping the predicted labels nearly unchanged (i.e., $\Delta\text{Label}$ is nearly zero) after perturbation.
 Answer to RQ2.
@@
 In other words, the predicted labels of many nodes have been changed after the attacks, and this violates our objective.
-5.3 Experiments to Answer RQ3: Attack Transferability
+VI-C Experiments to Answer RQ3: Attack Transferability
 Answer to RQ3.
-From Table 2, we can see that some post-hoc explainers, including Grad, GuidedBP, PGMExplainer, and IG, suffer from poor explanation accuracy on the original graphs, even giving lower accuracy than the Random Explainer.
+From Table III, we can see that some post-hoc explainers, including PGMExplainer, and IG, suffer from poor explanation accuracy on the original graphs, even giving lower accuracy than the Random Explainer.
 As a result, investigating the attack transferability on these explainers may not be meaningful.
-On the contrary, like PGExplainer, explainers such as GradCAM, GNNExplainer, and SubgraphX (gray) can achieve good performance on most original graphs.
+On the contrary, like PGExplainer, explainers such as GradCAM, GNNExplainer, SubgraphX, and GNN-LRP can achieve good performance on most original graphs.
 For those, the attacks optimized for PGExplainer are also harmful, leading to a dramatic degradation of explanation accuracy.
@@
 This demonstrates the transferability of the attacks generated by GXAttack.
-Other Analysis.
-We performed a sensitivity analysis for GXAttack in terms of maximal attack budget, maximal training epochs, and the loss trade-off hyperparameter.
-The results and analysis are deferred to Appendix I due to space reasons.
-Moreover, we also conducted some property analysis of GXAttack in terms of the prediction confidence, and node degree in Appendix H.
-Table 2: Transferability of attacks on selected synthetic datasets with a typical run.
-The results are consistent on other datasets and thus are deferred to Table 12 to save space.
+VI-D Sensitivity Analysis
+Specifically, we perform analysis in terms of $\Delta$GEA, $\Delta$Prob and used budget when varying the following hyperparameters, respectively: 1) the loss trade-off hyperparameter; 2) the maximally allowed perturbation budget (Max Budget) in Figure 2; and 3) the maximal training epochs of attacker in Figure 3.
+TABLE III: Transferability of attacks on synthetic datasets Syn1-6 with a typical run (Results on Syn7 are omitted due to excessive computation time).
 We report the explanation accuracy (i.e., GEA) on the original graph $\rightarrow$ explanation accuracy on the perturbed graph.
-Explainer Syn1 Syn2 Syn3 Syn4
-PGExplainer 0.728 $\rightarrow$ 0.510 0.648 $\rightarrow$ 0.347 0.485 $\rightarrow$ 0.304 0.471 $\rightarrow$ 0.311
-Grad 0.025 $\rightarrow$ 0.065 $0$rightarrow 0.001 $0$rightarrow 0.000 0.003 $\rightarrow$ 0.001
-GradCAM 0.745 $\rightarrow$ 0.551 0.708 $\rightarrow$ 0.397 0.510 $\rightarrow$ 0.300 0.511 $\rightarrow$ 0.310
-GuidedBP $0$rightarrow 0.000 $0$rightarrow 0.000 $0$rightarrow 0.000 $0$rightarrow 0.000
-IG 0.118 $\rightarrow$ 0.104 0.134 $\rightarrow$ 0.103 0.136 $\rightarrow$ 0.101 0.109 $\rightarrow$ 0.092
-GNNExplainer 0.738 $\rightarrow$ 0.550 0.711 $\rightarrow$ 0.397 0.539 $\rightarrow$ 0.341 0.514 $\rightarrow$ 0.327
-SubgraphX 0.113 $\rightarrow$ 0.015 0.532 $\rightarrow$ 0.375 0.296 $\rightarrow$ 0.141 0.417 $\rightarrow$ 0.282
-PGMExplainer 0.015 $\rightarrow$ 0.064 0.046 $\rightarrow$ 0.039 0.031 $\rightarrow$ 0.034 0.036 $\rightarrow$ 0.041
-RandomExplainer 0.184 $\rightarrow$ 0.146 0.170 $\rightarrow$ 0.096 0.135 $\rightarrow$ 0.102 0.124 $\rightarrow$ 0.090
-6 Related Work
-We only briefly revisit closely related work due to space constraints. See Appendix F for more detail.
-GNN Explanations.
-Yuan et al.
-(2022); Li et al.
-(2022); Kakkad et al.
-(2023) perform comprehensive surveys and propose excellent taxonomies of GNN explanation methods.
-Specifically, GNN explanations methods mainly include: 1) Gradients-based methods: Grad (Simonyan et al., 2013), GradCAM (Pope et al., 2019), GuidedBP (Baldassarre and Azizpour, 2019), Integrated Gradients (IG) (Sundararajan et al., 2017); 2) Decomposition-based methods: GraphLRP (Schwarzenberg et al., 2019), GNN-LRP (Schnake et al., 2021); 3) Surrogate-based methods: GraphLIME (Huang et al., 2022), PGM-Explainer (Vu and Thai, 2020); 4) Generation-based methods: XGNN (Yuan et al., 2020); 5) Perturbation-based methods: GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), GraphMask (Schlichtkrull et al., 2020), SubgraphX (Yuan et al., 2021); and 6) CF-based methods: CF-GNNExplainer (Lucic et al., 2022), RCExplainer (Bajaj et al., 2021).
-Robustness/Stability of GNN Explanations.
-Yuan et al.
-(2022); Agarwal et al.
-(2022) point out that a reliable GNN explainer should exhibit stability, namely minor perturbations that do not impact the model predictions should not largely change the explanations.
-Particularly, Agarwal et al.
-(2022) theoretically analyze the stability of three GNN explainers, including vanilla Grad, GraphMask and GraphLIME, for which they derive upper bounds on the instability of explanations.
-Moreover, Bajaj et al.
-(2021); Fang et al.
-(2024); Wang et al.
-(2024); Fan et al.
-(2023) are closely related but are different from our work (as compared in Section 1 and detailed in Appendix F).
-Adversarial Robustness of GNNs.
-Starting with the seminal works (Zügner et al., 2018; Dai et al., 2018), that both proposed adversarial attacks on GNNs, a rich literature formed in the realm of GNN-specific adversarial attacks, defenses, and certifications (Jin et al., 2021; Günnemann, 2022).
-Most attacks primarily focus on altering predictions via perturbations of the discrete graph structure (edge insertions/deletions), either between existing nodes (Zügner et al., 2018) or via the insertion of new (adversarial) nodes (Zou et al., 2021).
-Optimization-based attacks either operate globally (number of edge insertions/deletions) (Xu et al., 2019; Geisler et al., 2021), alter a local neighborhood (Zügner et al., 2018), or can handle a combination of both constraints (Gosch et al., 2023).
-While adversarial attacks on GNN explanation methods may benefit from optimization methods over the discrete graph structure, a careful discussion and derivation of appropriate attack objectives and metrics was missing.
-7 Conclusions
+Explainer Syn1 Syn2 Syn3 Syn4 Syn5 Syn6
+PGExplainer 0.728 $\rightarrow$ 0.510 0.648 $\rightarrow$ 0.347 0.485 $\rightarrow$ 0.304 0.471 $\rightarrow$ 0.311 0.450 $\rightarrow$ 0.285 0.332 $\rightarrow$ 0.225
+GradCAM 0.745 $\rightarrow$ 0.551 0.708 $\rightarrow$ 0.397 0.510 $\rightarrow$ 0.300 0.511 $\rightarrow$ 0.310 0.455 $\rightarrow$ 0.278 0.414 $\rightarrow$ 0.241
+IG 0.118 $\rightarrow$ 0.104 0.134 $\rightarrow$ 0.103 0.136 $\rightarrow$ 0.101 0.109 $\rightarrow$ 0.092 0.116 $\rightarrow$ 0.085 0.078 $\rightarrow$ 0.065
+GNNExplainer 0.738 $\rightarrow$ 0.550 0.711 $\rightarrow$ 0.397 0.539 $\rightarrow$ 0.341 0.514 $\rightarrow$ 0.327 0.483 $\rightarrow$ 0.283 0.431 $\rightarrow$ 0.205
+SubgraphX 0.113 $\rightarrow$ 0.015 0.532 $\rightarrow$ 0.375 0.296 $\rightarrow$ 0.141 0.417 $\rightarrow$ 0.282 0.394 $\rightarrow$ 0.312 0.408 $\rightarrow$ 0.305
+PGMExplainer 0.015 $\rightarrow$ 0.064 0.046 $\rightarrow$ 0.039 0.031 $\rightarrow$ 0.034 0.036 $\rightarrow$ 0.041 0.038 $\rightarrow$ 0.043 0.037 $\rightarrow$ 0.040
+GNN-LRP 0.752 $\rightarrow$ 0.573 0.713 $\rightarrow$ 0.394 0.509 $\rightarrow$ 0.300 0.515 $\rightarrow$ 0.311 0.457 $\rightarrow$ 0.284 0.410 $\rightarrow$ 0.241
+RandomExplainer 0.184 $\rightarrow$ 0.146 0.170 $\rightarrow$ 0.096 0.135 $\rightarrow$ 0.102 0.124 $\rightarrow$ 0.090 0.128 $\rightarrow$ 0.100 0.133 $\rightarrow$ 0.086
+Figure 2: Sensitivity analysis w.r.t. maximally allowed perturbation budget (Max Budget) on Syn2 using GXAttack.
+Trade-off Hyperparameter vs Attack Performance.
+Our sensitivity analysis shows that the attack performance (in term of $\Delta$GEA), $\Delta$Prob, and the used budget are nearly constant when we vary the trade-off hyper-parameter $\beta$’s value from $0.000001$ to 10.
+We omit the figures here.
+Attack Budget vs Attack Performance.
+From Figure 2, we can see that as the Max Budget increases from 1 to 10, $\Delta$GEA and $\Delta$Prob rise sharply, showing large gains from higher allowed budgets.
+Beyond a (maximally allowed) budget of 10, $\Delta$GEA and $\Delta$Prob both stabilize, indicating no further gains.
+The underlying reason is that the used budget stop increasing after 10 (maximally allowed) when we set the training epochs as 100.
+In Tables II and III, we report the results with a maximum budget equal to 15.
+Attack Training Epochs vs Attack Performance.
+From Figure 3, we observe a steady increase in the used budget as epochs increase.
+However $\Delta$Prob first increases and then decreases (due to overfitting) when increasing the training epochs.
+Similarly, $\Delta$GEA first increases and then decreases (due to overfitting) with the increase of training epochs.
+For computation time reason, we only report the results with 100 epochs in Tables II and III.
+However, it is important to note that the attack performance (in terms of $\Delta$GEA) can be largely improved if we increase the number of training epochs.
+For instance, $\Delta$GEA can be improved from 0.269 to 0.364 if we increase the training epochs from 100 to 300.
+Figure 3: Sensitivity analysis w.r.t. maximal training epochs on Syn2 using GXAttack.
+VI-E Property Analysis
+We also conducted some property analysis of GXAttack in terms of the prediction confidence, and node degree.
+Property Analysis on Prediction Confidence.
+From Figure 4(a), we can clearly see that nodes in ‘G9’ tend to have lower original explanation accuracy.
+As a result, as shown in Figure 4(b), the attack performance of GXAttack on these nodes (in ‘G9’) tend to be less successful (i.e., smaller $\Delta\text{GEA}$).
+This is because the nodes in ‘G9’ leave no much room for perturbations under the constraint that the predictions remain unchanged, namely much smaller probability changes as shown in Figure 4(c).
+Property Analysis on Node Degree.
+From Figure 5(a), we can see that nodes with higher degrees tend to have lower original explanation accuracy.
+Figure 5(b) shows that the attack performance of GXAttack on these nodes with higher degrees tends to be less successful (i.e., smaller $\Delta\text{GEA}$).
+Possible reasons are that 1) the original explanation accuracy tends to be lower; and 2) small perturbations on nodes with high degrees have limited impact on the predictions, resulting in smaller probability changes as shown in Figure 5(c).
+(a) Prediction confidence vs. original explanation accuracy on Syn2.
+(b) Prediction confidence vs. attack performance ($\Delta\text{GEA}$) on Syn2.
+(c) Prediction confidence vs. probability change after attacks on Syn2.
+Figure 4: Property analysis regarding prediction confidence on Syn2. Other datasets show consistent results and are omitted.
+(a) Node degree vs. original explanation accuracy on Syn2.
+(b) Node degree vs. attack performance ($\Delta\text{GEA}$) on Syn2.
+(c) Node degree vs. probability change after attacks on Syn2.
+Figure 5: Property analysis regarding node degree on Syn2. Other datasets show consistent results and thus are omitted.
+VII Conclusions
 GNN explanations for enhancing transparency and trust of graph neural networks are becoming increasingly important in decision-critical domains.
@@
 Future research should focus on developing explanation methods that are both interpretable and robust against adversarial attacks, as enhancing the stability of GNN explanations is crucial for ensuring their consistency and reliability in practical applications.
-The limitations and broader impact of our work are discussed in Appendices J and  K, respectively.
 Acknowledgements
 Zhong Li and Matthijs van Leeuwen: this publication is part of the project Digital Twin with project number P18-03 of the research programme TTW Perspective, which is (partly) financed by the Dutch Research Council (NWO).
-Simon Geisler: This research was supported by the Helmholtz Association under the joint research school “Munich School for Data Science – MUDS".
-Appendix for paper submission Explainable Graph Neural Networks Under Fire
-Appendix A Revisiting PGExplainer
-For completeness, we revisit PGExplainer as follows.
-For brevity, we denote $\mathcal{G}(v)$ as $\mathcal{G}$ (and $\mathcal{G}_{s}(v)$ as $\mathcal{G}_{s}$ analogously) without causing confusions throughout this subsection.
-Formalisation.
-Given a graph $\mathcal{G}$ and the target node $v$, to explain why $f_{\theta}(\mathcal{G}(v))=c$, PGExplainer will provide a local explanation (i.e.
-explanatory subgraph) $\mathcal{G}_{s}(v)=(\mathbf{A}_{s},\mathbf{X}_{s})$, where $\mathbf{A}_{s}$ is the subgraph structure and $\mathbf{X}_{s}$ is the node features associated with nodes residing in this subgraph.
-Specifically, PGExplainer maximises the mutual information between the GNN classifier’s prediction $\mathbf{y}$ and the explanatory subgraph $\mathcal{G}_{s}$:
-$\underset{\mathcal{G}_{s}}{\max}~{}\text{MI}(\mathbf{y},\mathcal{G}_{s}):=H(\mathbf{y})-H(\mathbf{y}|\mathcal{G}_{s})$ (17)
-where $H(\mathbf{y})$ is the entropy of $f_{\theta}(\mathcal{G})$, and $H(\mathbf{y}|\mathcal{G}_{s})$ is the conditional entropy of the GNN’s prediction on the explanatory graph, namely $f_{\theta}(\mathcal{G}_{s})$.
-As $\theta$ and $\mathcal{G}$ are fixed at the explanation stage (i.e.
-$\mathbf{y}$ is also fixed), maximising Eq 17 is equivalent to minimising $H(\mathbf{y}|\mathcal{G}_{s})$.
-Relaxations.
-However, directly minimising $H(\mathbf{y}|\mathcal{G}_{s})$ is intractable as there are $\mathcal{O}(2^{|\mathcal{E}|})$ candidates for $\mathcal{G}_{s}$.
-Therefore, they considered three relaxations: 1) they assume that the explanatory subgraph $\mathcal{G}_{s}$ is a Gilbert random graph, namely the selections of edges from $\mathcal{G}$ are conditionally independent to each other.
-As a result, the probability of graph $\mathcal{G}$ can be factorised as $P(\mathcal{G})=\prod_{(i,j)\in\mathcal{E}}P(e_{ij})$.
-Moreover, they instantiate $P(e_{ij})$ using a Bernoulli distribution $B(\lambda_{ij})$, namely $e_{ij}\sim B(\lambda_{ij})$ with $P(e_{ij}=1)=\lambda_{ij}$.
-As a result, Eq 17 is equivalent to:
-$\underset{\mathcal{G}_{s}}{\min}~{}H(\mathbf{y}|\mathcal{G}_{s})=\underset{\mathcal{G}_{s}}{\min}~{}\mathbb{E}_{\mathcal{G}_{s}}[H(\mathbf{y}|\mathcal{G}_{s})]\approx\underset{\Lambda}{\min}~{}\mathbb{E}_{\mathcal{G}_{s}\sim q(\Lambda)}[H(\mathbf{y}|\mathcal{G}_{s})],$ (18)
-where $q(\Lambda)$ describes the distribution of explanatory subgraph $\mathcal{G}_{s}$ that is parameterized by $\lambda$’s in $\Lambda$.
-2) they relax the edge weights from binary variables $\{0,1\}$ to continuous variable $[0,1]$ using reparameterization trick.
-As a result, they can utilise gradient-based methods to optimize the objective function.
-More specifically, the continuous edge weight $\hat{e}_{ij}$ is calculated as $\hat{e}_{ij}=sigmoid[(\ln(\epsilon)-\ln(1-\epsilon)+\omega_{ij})/\tau]$, where $\epsilon\sim\text{U}(0,1)$, $\tau$ is the temperature hyperparameter, and $\omega_{ij}$ is the learnable parameter in $\Omega$.
-Hence, Eq 18 is equivalent to:
-$\underset{\Omega}{\min}~{}\mathbb{E}_{\epsilon\sim\text{U}(0,1)}[H(\mathbf{y}|\mathcal{G}_{s})].$ (19)
-3) for efficiency, they replace conditional entropy $H(\mathbf{y}|\mathcal{G}_{s})$ with cross entropy $H(\mathbf{y},\mathbf{y}_{s})$, and further apply Monte Carlo to approximate it:
-$\underset{\Omega}{\min}~{}\frac{-1}{K}\sum_{k=1}^{K}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G})\ln(P_{\theta}(y=c|\mathcal{G}_{s}^{(k)})),$ (20)
-where $K$ is the number of sampled explanatory subgraph $\mathcal{G}_{s}^{(k)}$, and $\theta$ is the set of parameters in trained $f_{\theta}(\cdot)$.
-Generalisations.
-If we consider $\Omega$ in Eq 20 as independent variables, then we have to retrain it each time when we want to explain an unseen instance.
-To mitigate this issue, they employ a parameterized network to learn $\Omega$, namely $\Omega=h_{\Phi}(\mathcal{G},\mathbf{Z})$, where $\mathbf{Z}$ is the learned node embedding matrix before inputting to the downstream tasks (namely the result of component 1 in GNN classifier).
-Besides, $h(\cdot)$ is the explanation network that is shared by all edges, and $\Phi$ is the set of its parameters.
-As a result, in the collective setting Eq 20 becomes:
-$\underset{\Phi}{\min}~{}\sum_{i\in\mathcal{I}}\sum_{k=1}^{K}\sum_{c=1}^{C}P_{\theta}(y=c|\mathcal{G}^{(i)})\ln(P_{\theta}(y=c|\mathcal{G}_{s}^{(i,k)})),$ (21)
-where $\mathcal{G}^{(i)}$ is the $i$-th input graph from a set $\mathcal{I}$, and $\mathcal{G}_{s}^{(i,k)}$ is its $k$-th sampled explanatory subgraph.
-Finally, they implement the explanatory network $h(\cdot)$ as a MLP, which returns the edge importance as explanations.
-For node classification, the explanation is $\omega_{ij}=\text{MLP}_{\Phi}([\mathbf{z}_{i},\mathbf{z}_{j};\mathbf{z}_{v}])$ for node $v$.
-Appendix B Projected Gradient Descent
-After each gradient update $\tilde{\mathbf{P}}_{t}\in\mathbb{R}^{n\times n}$ although it must be in $[0,1]^{n\times n}$ for its interpretation as Bernoulli random variable.
+Simon Geisler: This research was supported by the Helmholtz Association under the joint research school “Munich School for Data Science – MUDS”.
+Appendix A: Projected Gradient Descent
+After each gradient update, $\tilde{\mathbf{P}}_{t}\in\mathbb{R}^{n\times n}$, although it must lie in $[0,1]^{n\times n}$ for its interpretation as a Bernoulli random variable.
 Moreover, we must ensure that sampling $\mathbf{P}_{t}\in\{0,1\}^{n\times n}\sim\text{Bernoulli}(\tilde{\mathbf{P}}_{t})$ will yield a discrete perturbation $\hat{\mathbf{A}}=\mathbf{A}\oplus\mathbf{P}_{t}$ obeying the admissible perturbations $\|\mathbf{A}-\hat{\mathbf{A}}\|_{0}<B$ with sufficient likelihood.
-Following Xu et al.
-(2019); Geisler et al.
-(2021), we ensure that $\mathbb{E}_{\mathbf{P}\sim\text{Bernoulli}(\tilde{\mathbf{P}_{t}})}[\|\mathbf{P}\|_{0}]\leq B$ and then obtain the final perturbation form a small set of samples.
+Following [46, 47], we ensure that $\mathbb{E}_{\mathbf{P}\sim\text{Bernoulli}(\tilde{\mathbf{P}_{t}})}[\|\mathbf{P}\|_{0}]\leq B$ and then obtain the final perturbation form a small set of samples.
 To guarantee that $\tilde{\mathbf{P}}_{t}$ parametrizes a valid distribution and that we obey the budget in expectation, we employ the projection $\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}}_{t})\leq B]}(\tilde{\mathbf{P}}_{t})$.
@@
 Thus, after each gradient update, we solve
-$\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}})\leq B]}(\tilde{\mathbf{P}})=\arg\min_{\tilde{\mathbf{S}}}\|\tilde{\mathbf{P}}-\tilde{\mathbf{S}}\|_{F}^{2}\qquad\text{subject to }\;$ $\sum_{i,j}\tilde{S}_{i,j}\leq B$ (22)
-$\;\tilde{\mathbf{S}}\in[0,1]^{n\times n}$
-for clipping to the hyperplane and projecting back towards the simplex using a bisection search. See Xu et al. (2019) for details and the derivation.
-Appendix C Complexity Analysis of the GXAttack Algorithm
-Algorithmic Components
-  1. 1.
-    Initial Predictions and Explanations:
-    • $f_{\theta}(\mathbf{A},\mathbf{X})$: Predicting using the GNN on the graph $\mathcal{G}$.
-The complexity depends on the GNN architecture but typically involves matrix multiplications between the adjacency matrix $\mathbf{A}$ and the feature matrix $\mathbf{X}$.
-    • $g_{\lambda}(f_{\theta}(\mathbf{A},\mathbf{X}))$: Generating explanations based on the GNN predictions.
-The complexity is generally proportional to that of the GNN prediction if the explainer works by perturbing inputs and observing changes in output.
-  2. 2.
-    Loop for Gradient Updates:
-    • Prediction and Explanation on Perturbed Graph: This operation is similar in complexity to the initial prediction and explanation but is repeated $T$ times (number of epochs).
-    • Gradient Computation and Update:
-      • Calculating the gradient $\nabla_{\tilde{\mathbf{P}}_{t-1}}\mathcal{L}$ involves computing derivatives of the loss function with respect to the perturbation matrix $\tilde{\mathbf{P}}$.
-This can be expensive as it depends on the sensitivity of both the GNN predictions and explanations to changes in the adjacency matrix.
-      • Updating $\tilde{\mathbf{P}}_{t}$ involves a simple addition and a projection operation.
-The projection step is constrained by the perturbation budget $B$, but the actual operation is typically linear with respect to the number of potential edges $n^{2}$.
-  3. 3.
-    Binary Perturbation Sampling:
-    • Sampling from a Bernoulli distribution for each potential edge (pair of nodes), which has a complexity proportional to $n^{2}$.
-Computational Complexity
-We discuss the dense case here and refer to the summary below for a discussion of GNNs/explainers implemented with sparse matrices.
-  1. 1.
-    Initial Operations: $O(f_{\theta}+g_{\lambda})$, typically $O(n^{2})$ for a dense GNN and explainer (sometimes even $O(n^{3})$).
-  2. 2.
-    Each Epoch:
-    • Prediction and explanation on perturbed graph: $O(2T\cdot(f_{\theta}+g_{\lambda}))$.
-    • Gradient computation and update: Dependent on the efficiency of backpropagation through the GNN, but generally $O(T\cdot n^{2})$ for gradient computation and a similar order for updates.
-  3. 3.
-    Sampling Step: $O(n^{2})$.
-Space Complexity
-  1. 1.
-    Space for Storing Matrices: Storing $\mathbf{A}$, $\mathbf{X}$, $\tilde{\mathbf{P}}$ each requires $O(n^{2})$ space.
-  2. 2.
-    Additional Space for Operations: Space required for intermediate gradients and outputs during predictions, typically $O(n^{2})$.
-Summary
-Since the explainer uses dense matrices, we also choose an attack operating on all $n^{2}$ edges.
-Thus, the total complexity $O((T+1)\cdot(f_{\theta}+g_{\lambda})+n^{2})$ since we have $T$ attack steps plus the initial predictions as well as explanations and, last, we sample the final perturbations.
-Assuming that $O(f_{\theta})=O(g_{\lambda})=O(n^{2}d+nd^{2})$ and $d\ll n$, this yields a total time complexity of $O(n^{2})$ with number of attack iterations $T\ll n$, and number of nodes $n$.
-This is the same asymptotic complexity as the dense GNN $f_{\theta}$ and explainer $g_{\lambda}$.
-Assuming GNN and explainer are in $O(f_{\theta})=O(g_{\lambda})=O(Ed+nd^{2})$ with number of edges $E$, a dense attack would dominate the overall complexity ($O(n^{2})$).
-An immediate remedy would yield PRBCD (Geisler et al., 2021) since its complexity is $O(B+E)$ with edge flipping budget $B$ and all additional components (e.g., loss function) could be implemented in $O(E)$ with sparse matrices.
-This yields an overall time and space complexity of $O(E)$, assuming $B\ll E$ and hidden dimensions $d\ll n$, which is the same asymptotic complexity as the sparse GNN $f_{\theta}$ and explainer $g_{\lambda}$.
-Appendix D Datasets Description
-Overall, we consider the following synthetic datasets, which are summarised in Table 3: 1) synthetic datasets with different size via varying the following hyper-parameters in ShapeGGen: num_subgraphs for the number of subgraphs and subgraph_size for the number of expected nodes in each subgraph; 2) synthetic datasets with different types of explanatory graphs: shape as “circle" or “house"; and 3) synthetic datasets with different node degrees: prob_connection with different values.
-Table 3: Summary of synthetic datasets with ground-truth explanations.
-#Nodes and #Edges represent the number of expected nodes and edges, respectively.
-We do not consider larger datasets for two reasons: 1) explanation accuracy will be very low ($<0.3$) for most explainers, and it is not meaningful to attack low accuracy explainers; 2) computation will be slow for the attacker (and the explainer).
-Dataset Syn1 Syn2 Syn3 Syn4 Syn5 Syn6 Syn7
-Shape House House House Circle House House House
-#Subgraphs 10 50 50 50 50 50 100
-Subgraph Size 6 6 10 10 10 10 10
-P(Connection) 0.3 0.06 0.06 0.06 0.12 0.20 0.06
-#Classes 2 2 2 2 2 2 2
-#Nodes 59 299 521 511 489 492 1018
-#Edges 164 970 1432 1314 1664 1898 3374
-Average Degree 2.8 3.2 2.7 2.6 3.4 3.9 3.3
-Max Budget 10 15 15 15 15 15 15
-Appendix E Presumptions Checking Results
-Table 4: Complete results on Syn1.
+$\prod_{\mathbb{E}[\text{Bernoulli}(\tilde{\mathbf{P}})\leq B]}(\tilde{\mathbf{P}})=\arg\min_{\tilde{\mathbf{S}}}\|\tilde{\mathbf{P}}-\tilde{\mathbf{S}}\|_{F}^{2}$ (21)
+subject to $\qquad\text{subject to }\;\sum_{i,j}\tilde{S}_{i,j}\leq B~{}\text{and}~{}\tilde{\mathbf{S}}\in[0,1]^{n\times n}$
+for clipping to the hyperplane and projecting back towards the simplex using a bisection search. See [46] for details and the derivation.
+Appendix B: Presumptions Checking Results
+(a) Syn1
+(b) Syn2
+(c) Syn3
+(d) Syn4
+(e) Syn5
+(f) Syn6
+(g) Syn7
+Figure 6: Comparative analysis of explanation accuracy with respect to prediction correctness on all synthetic datasets.
+The results show that explanation accuracy is not necessarily correlated with the prediction correctness.
+Particularly, we did not observe that wrongly predicted nodes tend to have lower explanation accuracy.
+TABLE IV: Complete results on Syn1-Syn7.
 O.
@@
 Besides, $\Delta$Label is the average change of predicted labels, while $\Delta$Prob denotes the average absolute change of predicted probability (of the original predicted class).
-“G5" is the set of nodes of which the original prediction confidence located in [0.5,0.6).
-Similarly, “G6" for [0.6,0.7), “G7" for [0.7,0.8), “G8" for [0.8,0.9), “G9" for [0.9,1.0], and “All" for all nodes.
-Confidence Group G5 G6 G7 G8 G9 All
-O. GEA 0.696 0.763 0.726 0.775 0.683 0.728
+“G5” is the set of nodes of which the original prediction confidence located in [0.5,0.6).
+Similarly, “G6” for [0.6,0.7), “G7” for [0.7,0.8), “G8” for [0.8,0.9), “G9” for [0.9,1.0], and “All” for all nodes.
+Dataset Confidence Group G5 G6 G7 G8 G9 All
+Syn1 O. GEA 0.696 0.763 0.726 0.775 0.683 0.728
 P. GEA ()($\downarrow$)( ↓ ) 0.403 0.463 0.485 0.584 0.679 0.560
@@
 Simcos ()($\downarrow$)( ↓ ) 0.817 0.868 0.898 0.937 0.987 0.923
-#Pert ()($\downarrow$)( ↓ ) 6.5 5.0 4.5 2.7 0.2 3.0
+#Pert. ()($\downarrow$)( ↓ ) 6.5 5.0 4.5 2.7 0.2 3.0
 $\Delta$Label ()($\downarrow$)( ↓ ) 0 0 0 0 0 0
 $\Delta$Prob ()($\downarrow$)( ↓ ) 0.327 0.204 0.147 0.065 0.002 0.108
-Table 5: Complete results on Syn2. See Table 4 for the meaning of row names and column names.
-Confidence Group G5 G6 G7 G8 G9 All
-O. GEA 0.694 0.739 0.690 0.694 0.546 0.648
+Syn2 O. GEA 0.694 0.739 0.690 0.694 0.546 0.648
 P. GEA ()($\downarrow$)( ↓ ) 0.329 0.396 0.390 0.387 0.381 0.347
@@
 $\Delta$Prob ()($\downarrow$)( ↓ ) 0.353 0.268 0.181 0.103 0.003 0.143
-(a) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn1.
-(b) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn2.
-(c) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn3.
-(d) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn4.
-(e) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn5.
-(f) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn6.
-(g) Distribution of explanation accuracy w.r.t correctly and wrongly predicted nodes on Syn7.
-Figure 2: Comparative analysis of explanation accuracy with respect to prediction correctness on all synthetic datasets.
-The results show that explanation accuracy is not necessarily correlated with the prediction correctness.
-Particularly, we did not observe that wrongly predicted nodes tend to have lower explanation accuracy.
-Table 6: Complete results on Syn3. See Table 4 for the meaning of row names and column names.
-Confidence Group G5 G6 G7 G8 G9 All
-O. GEA 0.484 0.496 0.479 0.510 0.440 0.480
+Syn3 O. GEA 0.484 0.496 0.479 0.510 0.440 0.480
 P. GEA ()($\downarrow$)( ↓ ) 0.305 0.303 0.279 0.335 0.310 0.306
@@
 $\Delta$Prob ()($\downarrow$)( ↓ ) 0.242 0.206 0.166 0.099 0.039 0.165
-Table 7: Complete results on Syn4. See Table 4 for the meaning of row names and column names.
-Confidence Group G5 G6 G7 G8 G9 All
-O. GEA()($\downarrow$)( ↓ ) 0.500 0.500 0.493 0.495 0.416 0.471
+Syn4 O. GEA 0.500 0.500 0.493 0.495 0.416 0.471
 P. GEA ()($\downarrow$)( ↓ ) 0.309 0.314 0.319 0.326 0.296 0.311
@@
 $\Delta$Prob ()($\downarrow$)( ↓ ) 0.333 0.267 0.177 0.111 0.030 0.149
-Table 8: Complete results on Syn5. See Table 4 for the meaning of row names and column names.
-Confidence Group G5 G6 G7 G8 G9 All
-O. GEA 0.456 0.470 0.443 0.461 0.430 0.450
+Syn5 O. GEA 0.456 0.470 0.443 0.461 0.430 0.450
 P. GEA ()($\downarrow$)( ↓ ) 0.252 0.264 0.282 0.305 0.336 0.293
@@
 $\Delta$Prob ()($\downarrow$)( ↓ ) 0.287 0.226 0.147 0.085 0.031 0.138
-Table 9: Complete results on Syn6. See Table 4 for the meaning of row names and column names.
-Confidence Group G5 G6 G7 G8 G9 All
-O. GEA 0.415 0.412 0.407 0.382 0.291 0.332
+Syn6 O. GEA 0.415 0.412 0.407 0.382 0.291 0.332
 P. GEA ()($\downarrow$)( ↓ ) 0.260 0.249 0.239 0.259 0.217 0.231
@@
 $\Delta$Prob ()($\downarrow$)( ↓ ) 0.348 0.270 0.189 0.106 0.029 0.092
-Table 10: Complete results on Syn7. See Table 4 for the meaning of row names and column names.
-Confidence Group G5 G6 G7 G8 G9 All
-O. GEA 0.428 0.415 0.406 0.405 0.401 0.409
+Syn7 O. GEA 0.428 0.415 0.406 0.405 0.401 0.409
 P. GEA ()($\downarrow$)( ↓ ) 0.266 0.250 0.294 0.263 0.278 0.263
@@
 $\Delta$Prob ()($\downarrow$)( ↓ ) 0.319 0.253 0.169 0.097 0.035 0.147
-(a) Explanation accuracy changes after attacks for nodes with prediction confidence ($0.60.5\leq p<$ 0.6).
-One can see that the attacks tend to be more successful if the original explanation accuracy is higher.
-(b) Explanation accuracy changes after attacks for nodes with different prediction confidence.
-Figure 3: Comparison of explanation accuracy changes after attacks across different prediction confidences on Syn2.
-The starting point of each directed line segment represents the original accuracy, the end point (i.e., the arrow position) represents the accuracy after perturbation, and the length of the line segment represents the change in explanation accuracy.
-Appendix F Related Work in More Detail
-Related work includes adversarial attacks and defenses on graphs, GNN explanations, attacks and defenses of traditional XAI methods, and stability of GNN explanation methods.
-F.1 Adversarial Attacks and Defenses on Graphs
-Graph adversarial attacks attempt to deliberately manipulate or deceive graph learning models by making small, often imperceptible/unnoticeable perturbations to the input graph’s structure (i.e., edge/node additions/deletions) and/or attributes (i.e., attribute perturbations).
-Particularly, most existing adversarial attacks and defenses methods focus on node-level operations and tasks Jin et al.
-(2021).
-Furthermore, these methods can be roughly categorised into: 1) white-box attacks, 2) gray-box attacks, and 3) black-box attacks according to the attacker’s knowledge.
-Specifically, white-box attacks usually assume that attackers can calculate gradient via model parameters, which are generally not feasible in real-world applications.
-Meanwhile, gray box attacks usually assume that they have access to the training data and corresponding labels (node-level or graph-level).
-On this basis they usually train a surrogate model to mimic the real model (not accessible), and then attack the surrogate model.
-Black box attacks can only access the adjacency matrix, node attribute matrix, and the output of the victim GNN model (model parameters, ground-truth labels and predicted probabilities are not available).
-Most black-box attacks are constructed based on the attack transferability assumption (namely attacks are effective cross different models) and perform queries from victim model.
-F.2 GNN Explanations
-Yuan et al.
-(2022); Li et al.
-(2022); Kakkad et al.
-(2023) perform comprehensive surveys and propose excellent taxonomies of GNN explanation methods.
-GNN explanations methods can be divided into the following categories based on Yuan et al.
-(2022):
-  • Gradients-based methods: Vanilla Grad (Simonyan et al., 2013), GradCAM (Pope et al., 2019), GuidedBP (Baldassarre and Azizpour, 2019), Integrated Gradients (IG) (Sundararajan et al., 2017);
-  • Decomposition-based methods: GraphLRP (Schwarzenberg et al., 2019), Excitation BP, GNN-LRP (Schnake et al., 2021);
-  • Surrogate-based methods: GraphLIME (Huang et al., 2022), PGM-Explainer (Vu and Thai, 2020);
-  • Generation-based methods: XGNN (Yuan et al., 2020);
-  • Perturbation-based methods: GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), GraphMask (Schlichtkrull et al., 2020), SubgraphX (Yuan et al., 2021);
-  • CF-based methods: CF-GNNExplainer (Lucic et al., 2022), RCExplainer (Bajaj et al., 2021).
-Particularly, they are summarised in Table 11.
-Table 11: Summary of GNN explanation methods.
-Method Venue Datasets Metrics Code
-GNNExplainer NIPS’19 Synthetic (BA-Shapes, BA-Community, Tree-Cycles, Tree-Grid) for node classification; Mutag and Reddit-Binary for graph classification Explanation Accuracy Github
-GraphLIME TKDE’22 Cora, PubMed for node classification Noisy features ratio Github
-XGNN KDD’20 Synthetic (Is_Acyclic) and Mutag for graph classification Visualisation Github
-PGExplainer NIPS’20 Synthetic (BA-Shapes, BA-Community, Tree-Cycles, Tree-Grid) for node classification; Synthetic (BA-2motifs) and Mutag for graph classification Explanation Accuracy; Visualisation Github
-PGM-Explainer NIPS’20 Synthetic, Bitcoin-Alpha and Bitcoin-OTC for node classification; SuperPixel-Graph for graph classification Explanation Accuracy; Human-subjective test Github
-Gradients CVPR’19 Visual Genome scene-graphs, BBBP, BACE, and TOX21 for graph classification Fidelity, Contrastivity, Sparsity Github
-SubgraphX ICML’21 Synthetic (BA-Shapes) for node classification; Synthetic (BA-2Motifs), Mutag, BBBP, Graph-SST2 for graph classification Fidelity, Sparsity, Visualisation Github
-CF-GNNExplainer AIST’22 Synthetic (BA-Shapes, Tree-Cycles, Tree-Grid) for node classification; Fidelity, Sparsity, Explanation Size, Accuracy Github
-RCExplainer NIPS’21 Synthetic (BA-Shapes, BA-Community, Tree-Cycles, Tree-Grid) for node classification; Synthetic (BA-2motifs), Mutag, NCI1 for graph classification; Fidelity, Robustness, Accuracy, Efficiency Github
-V-InFoR NIPS’23 Synthetic (BA-3motifs), Mutag, Ogbg-molhiv, Ogbg-ppa for graph classification; Sufficient Accuracy, Necessary Accuracy, $F_{NS}$ Score Github
-GraphSVX ECML’21 Synthetic (BA-Shapes, BA-Community, Tree-Cycles, Tree-Grid) for node classification; Synthetic (BA-2motifs) and Mutag for graph classification; Accuracy, Robustness, Certainty Github
-F.3 Attacks and Defenses of Traditional XAI Methods
-It has been shown that traditional XAI methods (namely those designed to explain deep neural networks for image and text data) are susceptible to malicious perturbations (Heo et al., 2019; Zhang et al., 2020; Ghorbani et al., 2019).
-Please refer to Baniecki and Biecek (2024) for a comprehensive survey.
-F.4 Robustness of GNN Explanations
-OAR.
-Fang et al.
-(2024) present a novel metric for evaluating GNN explanations, termed Out-Of-Distribution-resistant Adversarial Robustness (OAR).
-OAR aims to solve the limitations of current removal- and generative-based evaluations.
-More concretely, they evaluate post-hoc explanation subgraphs by computing their robustness under attack, which consists of three steps: 1) they formulate the adversarial robustness tailored for GNN explanations problem, which is the minimum adversarial perturbation on the structure of complementary subgraph (i.e., removing explanatory subgraph from the original graph); In other words, a good explanation subgraph will yield high “such-defined” adversarial robustness; 2) they introduce a tractable and easy-to-implement objective of adversarial robustness for GNN explanations; after relaxation, the objective becomes a constrained graph generation problem (namely random perturbations) rather than an adversarial attack problem; 3) they present an OOD reweighting block (which is instantiated using a graph anomaly detector VGAE) that aims to confine the evaluation on original data distribution.
-The OOD score is used as the weight of each prediction when computing the expectation of generated graph’s prediction.
-They evaluate the performance of post-hoc explainers including GNNExplainer, PGExplainer, Refine, SA, CXPlain, and GradCAM, on BA3, TR3, Mutag and MNIST datasets for graph classification task.
-They consider ground-truth explanations, consistency with metric that is defined using ground-truth explanations, and consistency with human intuition.
-This work is an initial attempt to explore evaluation metrics for GNN explanations from the perspectives of adversarial robustness and resistance to OOD.
-V-InFoR.
-(Wang et al., 2024) point out that “existing GNN explainers are not robust to the structurally corrupted graphs, namely graphs with noisy or adversarial edges." This study for the first time investigates the negative effect of minor structural corruptions (which do not change the predictions) and severe structural corruptions (which change the predictions) on GNN explainers.
-Particularly, they provide quantitative evidence that existing GNN explainers are fragile to structurally corrupted graphs.
-They evaluate 6 GNN explainers (GradCAM, IG, PGM-Explainer, GNNExplainer, PGEXplainer, ReFine) on BA-3Motifs dataset under two settings: 1) minor corruptions where they select the perturbed graphs whose predictions are not changed after randomly flipping 20% edges; 2) severe corruptions where they employ adversarial attack algorithm GRABNEL to perturb at most 10% edges to change the predictions.
-Their results shows that minor corruptions lead to an average of 31.15% explanation performance degradation, and severe corruptions lead to 40.64%.
-Particularly, V-InfoR consists of two modules: 1) a robust graph representation learner that aims to learn a variational distribution that contains common information shared a large amount of graphs; 2) an adaptive explanation generator that samples explanatory subgraph from the variational distribution.
-RCExplainer.
-Bajaj et al.
-(2021) aim to find a small set of edges of the input graph such that the prediction result will substantially change if we remove those edges.
-Specifically, they propose RCExplainer to generate robust counter-factual explanations for GNNs in two steps: 1) they employ a set of decision regions to model the decision logic of a GNN, where each decision region (corresponding to a class) determines the predictions on multiple graphs that are predicted to be the same class; 2) they leverage a DNN model to explore the decision logic (namely the linear decision boundaries of the decision regions), and thereby extract robust counter-factual explanations as a small set of edges of the input graph.
-Particularly, the linear decision boundaries of the decision region (of a specific class) capture the common decision logic on all graphs inside the decision region, making it not easy to overfit the noise of some individual graphs.
-This method is only applicable to GNNs that belong to Piecewise Linear Neural Networks (e.g., ReLU, Leaky ReLU, and Maxout).
-They point out (for the first time) that most methods such as PGExplainer, Gradients, Graph Attentions, and GNNExplainer are not robust to noise, namely slightly random perturbations on input graph may cause that the generated explanations substantially change while the predicted labels keep unchanged.
-(They argue that this is because the explanation of every individual prediction is independently optimized to maximize its correlation with the prediction, and thus can easily cause overfitting to noise in the data.) Moreover, they indicate that these explanation methods are usually not counter-factual: “removing the explanation subgraph from the input graph does not necessarily change the prediction result due to the non-convexity of GNNs." However, they did not systematically investigate this phenomena.
-GEAttack.
-Fan et al.
-(2023) empirically demonstrate that GNN explanation methods can be employed as tools to detect the adversarial perturbations on graphs.
-On this basis, they propose an attack framework dubbed GEAttack that can jointly attack both GNN and its explanations.
-Specifically, they formulate GEAttack as a bi-level optimization problem: 1) they mimic the GNN explanation method optimization process to obtain graph explanations in the inner loop; 2) they compute the gradient of the attack objective w.r.t.
-the explanations in the outer loop.
-Moreover, they considered two GNN explanation methods, including GNNExplainer and PGExplainer.
-Appendix G Training Protocols and Compute Resources
-G.1 GNN Predictor Training Protocols
-To avoid pitfalls caused by weak GNN predictor Faber et al.
-(2021), we should train the GNN predictor close to its maximum possible performance.
-Moreover, when training the GNN predictor, we employ an Adam optimizer where the learning rate is 1e-2, weight decay is 1e-5, the number of epochs is 300, and the hidden dimension is 32.
-G.2 GNN Explainer Training Protocols
-We followed the authors’ recommendations for setting the hyperparameters of GNN explanation methods.
-We select top-$k~{}(k=25\%)$ important edges (also top-$k$ nodes and node features if necessary) to generate explanations for all GNN explainers.
-Particularly, when training the PGExplainer, the number of training epochs is 10.
-G.3 GNN Explanation Attacker Training Protocols
-We set the trade-off hyperparameter in the final loss as $0.1$ such that the magnitudes of the two loss terms are approximately the same.
-Besides, the learning rate in gradient ascent is set to $0.1$.
-Importantly, we set the maximal training epochs as 100 due to computation time reason, while further increasing this number can largely improve the attack performance (in terms of $\Delta$GEA).
-Moreover, we set the perturbation budget to 15 at maximum for synthetic data (note that we attempt to employ a hard minimum budget of 5 with 2000 maximum trials of Bernoulli sampling).
-We perform sensitivity analysis regarding these hyperparameters, and the results and analysis are detailed in Appendix I.
-G.4 Explanation Accuracy Computation
-The $get\_acc(\cdot)$ function of GraphXAI considers all non-zero edges as predicted positives (namely with binary code ‘1’) by default when computing JAC.
-Following Agarwal et al.
-(2023), we set the top 25% (of edges with the highest edge importance scores) as 1 and the rest as zeros in our evaluations when computing both GEA and cosine similarity.
-G.5 Software and Hardware
-The evaluated GNN explainers are implemented mainly based on GraphXAI  [^1] https://github.com/mims-harvard/GraphXAI (with MIT License)(Agarwal et al., 2023).
-Algorithms are implemented in Python 3.8 (using PyTorch Paszke et al.
-(2019) and PyTorch Geometric Fey and Lenssen (2019) libraries when applicable) and ran on internal clusters with AMD 128 EPYC7702 CPUs (with 512GB RAM).
-All experiments can be finished within 7 days if using 10 such machine instances.
-All code and datasets will be available on GitHub [^2] https://github.com/ZhongLIFR/GXAttack.
-Appendix H Additional Results
-H.1 Property Analysis Regarding Prediction Confidence
-(a) Prediction confidence VS original explanation accuracy on Syn2.
-(b) Prediction confidence VS attack performance (i.e. $\Delta\text{GEA}$) on Syn2.
-(c) Prediction confidence VS probability change after attacks on Syn2.
-Figure 4: Property analysis regarding prediction confidence on Syn2.
-Other datasets show consistent results and thus are omitted.
-From Figure 4(a), we can clearly see that nodes in ‘G9’ tend to have lower original explanation accuracy.
-As a result, as shown in Figure 4(b), the attack performance of GXAttack on these nodes (in ‘G9’) tend to be less successful (i.e., smaller $\Delta\text{GEA}$).
-This is because the nodes in ‘G9’ leave no much room for perturbations under the constraint that the predictions remain unchanged, namely much smaller probability changes as shown in Figure 4(c).
-H.2 Property Analysis Regarding Node Degree
-(a) Node degree VS original explanation accuracy on Syn2.
-(b) Node degree VS attack performance (i.e. $\Delta\text{GEA}$) on Syn2.
-(c) Node degree VS probability change after attacks on Syn2.
-Figure 5: Property analysis regarding node degree on Syn2.
-Other datasets show consistent results and thus are omitted.
-From Figure 5(a), we can see that nodes with higher degree tend to have lower original explanation accuracy.
-Figure 5(b) shows that the attack performance of GXAttack on these nodes with higher degree tend to be less successful (i.e., smaller $\Delta\text{GEA}$).
-Possible reasons are that 1) the original explanation accuracy tends to be lower; and 2) small perturbations on nodes with high degree have limited impact on the predictions, namely resulting small probability changes as shown in Figure 5(c).
-H.3 Attack Transferability Results
-Table 12: Transferability of attacks on synthetic datasets.
-We report the explanation accuracy on original graph $\rightarrow$ explanation accuracy on perturbed graph.
-The results on Syn7 are omitted due to the fact that runtime exceeded a 10-day limit for a single trial.
-Explainer Syn5 Syn6
-PGExplainer 0.450 $\rightarrow$ 0.285 0.332 $\rightarrow$ 0.225
-Grad $0$rightarrow 0.000 $0$rightarrow 0.000
-GradCAM 0.455 $\rightarrow$ 0.278 0.414 $\rightarrow$ 0.241
-GuidedBP $0$rightarrow 0.000 $0$rightarrow 0.000
-IG 0.116 $\rightarrow$ 0.085 0.078 $\rightarrow$ 0.065
-GNNExplainer 0.483 $\rightarrow$ 0.283 0.431 $\rightarrow$ 0.205
-SubgraphX 0.394 $\rightarrow$ 0.312 0.408 $\rightarrow$ 0.305
-PGMExplainer 0.038 $\rightarrow$ 0.043 0.037 $\rightarrow$ 0.040
-RandomExplainer 0.128 $\rightarrow$ 0.100 0.133 $\rightarrow$ 0.086
-Appendix I Sensitivity Analysis
-We provide the results of sensitivity analysis in this section.
-Specifically, we perform analysis in terms of $\Delta$GEA, $\Delta$Prob and used budget when varying the following hyperparameters, respectively: 1) the loss trade-off hyperparameter; 2) the maximally allowed perturbation budget (Max Budget) in Figure 6; and 3) the maximal training epochs of attacker in Figure 7;
-Trade-off Hyperparameter vs Attack Performance.
-Our sensitivity analysis shows that the attack performance (in term of $\Delta$GEA), $\Delta$Prob, and the used budget are nearly constant when we vary the trade-off hyper-parameter $\beta$’s value from $0.000001$ to 10.
-We omit the figures here.
-Attack Budget vs Attack Performance.
-Figure 6: Sensitivity analysis w.r.t.
-maximally allowed perturbation budget (Max Budget) on Syn2 using GXAttack.
-We can see that as the Max Budget increases from 1 to 10, $\Delta$GEA and $\Delta$Prob rise sharply, showing large gains from higher allowed budgets.
-Beyond a (maximally allowed) budget of 10, $\Delta$GEA and $\Delta$Prob both stabilize, indicating no further gains.
-The underlying reason is that the used budget stop increasing after 10 (maximally allowed) when we set the training epochs as 100.
-In Tables 1 and 2, we report the results with a maximum budget equal to 15.
-Attack Training Epochs vs Attack Performance.
-Figure 7: Sensitivity analysis w.r.t.
-maximal training epochs on Syn2 using GXAttack.
-We observe a steady increase in the used budget as epochs increase.
-However $\Delta$Prob first increases and then decreases when increasing the training epochs.
-Similarly, $\Delta$GEA first increases and then decreases with the increase of training epochs.
-For computation time reason, we only report the results with 100 epochs in Tables 1 and 2.
-However, it is important to note that the attack performance (in terms of $\Delta$GEA) can be largely improved if we increase the number of training epochs.
-For instance, $\Delta$GEA can be improved from 0.269 to 0.364 if we increase the training epochs from 100 to 300.
-Appendix J Limitations
-Our current work suffers from the the following limitations and we plan to mitigate them in the future:
-  • GXAttack suffers from low scalability (large requirement of memory), which however can be mitigated by using the Projected Randomized Block Coordinate Descent (PR-BCD) (Geisler et al., 2021);
-  • We only employed synthetic datasets.
-While real-world datasets can provide quantitative evaluation results using metrics like cosine similarity (to measure the consistency of explanations before and after attacks), these metrics may not accurately reflect the true effectiveness of the attacks (i.e., in terms of $\Delta$GEA).
-This discrepancy suggests that quantitative results can be challenging to interpret, and relying solely on such metrics might not yield meaningful insights.
-On the other hand, qualitative evaluations are also possible, but as machine learning researchers, our expertise lies not in the realm of such interpretations;
-  • GXAttack requires that the target GNN explainer to be differentiable, which is not always the case; However, the attack transferability shows its potential effectiveness on other GNN explainers;
-  • We (implicitly) impose an assumption that a small $L_{0}$ perturbation that is prediction-preserving also preserves the causal reasoning for the GNN prediction.
-Unlike images or texts, it is not easy to verify for graph data.
-However, we can leverage visualisations to help understand the reasoning process: Figures 8 and  9 give an example which shows the edge perturbations and the corresponding explanation changes after the attacks of GXAttack.
-Figure 8: Comparison of original and perturbed graph structures for the target node ‘19’ on graph Syn1.
-The explanation accuracy decreases from $0.818$ to $0.539$ while the prediction probabilities change from $[0.676,0.324]$ to $[0.844,0.156]$.
-Figure 9: The explanation matrices of node ‘19’ before and after attacks, and the difference matrix (original - perturbed).
-Appendix K Broader Impacts
-Our findings in this paper have significant implications for the development and deployment of GNNs in a variety of domains, particularly in decision-critical applications.
-Our research highlights a fundamental vulnerability in current post-hoc explanation methods for GNNs, emphasizing that these methods can be unreliable when subject to adversarial perturbations.
-This revelation calls for a reassessment of how explanations of GNN decisions are generated and trusted, particularly in sectors where transparency and accountability are paramount, such as healthcare, finance, and autonomous systems.
-By introducing the GXAttack, an optimization-based adversarial attack specifically designed to target GNN explanations, this paper sets a new standard in the evaluation of explanation methods.
-This approach not only sheds light on the robustness of existing techniques but also serves as a catalyst for the development of more resilient GNN explanation methods.
-Ensuring the reliability of explanations in adversarial environments is crucial, as stakeholders rely on these explanations to make informed decisions and justify the use of GNNs in sensitive and impactful contexts.
-Moreover, by making our implementation publicly available, we enhance transparency and facilitate widespread involvement in evaluating and enhancing the robustness of GNN explainers.
-This openness is anticipated to expedite progress in the field of research dedicated to developing more reliable and interpretable graph learning models.
-Such advancements are likely to increase trust and acceptance of GNN technologies.
-The primary aim of our research is to contribute to the development of safer and more comprehensible AI systems, ensuring their ethical and responsible application in critical domains.